### v1

# Thorough Analysis of $\pi^{-}p\rightarrow K^{*}\Sigma$ Reaction: Probing Nucleon Excited State and Constituent Scaling Laws

Abstract
In this study, applying the effective Lagrangian approach, we investigate the reaction $\pi^{-}p\rightarrow K^{*}\Sigma$ to explore the production of the nucleon excited state $N^{*}(2080)$ and the constituent field scaling laws, also known as the constituent counting rules. Specifically, we analyze two scenarios, where the $J^{P}$ quantum numbers of $N^{*}(2080)$ are $3/2^{-}$ and $1/2^{-}$. Our results show that the contribution from $s$-channel $N^{*}$ exchange dominates the total cross section, surpassing those from $t$-channel $K^{(*)}$ exchange and $u$-channel $\Sigma$ exchange. Additionally, we present predictions for the differential cross section of the $\pi^{-}p\rightarrow K^{*}\Sigma$ process. Notably, in the large momentum transfer region ($\theta=90^{\circ}$), the differential cross section in the $t$-distribution aligns with the energy-dependent scaling formula predicted by the constituent counting rule, specifically, $d\sigma/dt\sim 1/s^{8}$ for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction. Finally, we extend the 2-to-2 scattering process to the 2-to-3 Dalitz process and provide theoretical predictions for this scenario. Our findings suggest that the cross section for the $\pi^{-}p\rightarrow K^{*}\Sigma\rightarrow K^{+}\pi^{-}\Sigma^{0}$ process can reach several tens of $\mu b$ near the threshold, making it highly favorable for experimental measurement. However, the current scarcity of experimental data for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction at threshold energy limits our ability to precisely determine the properties of the nucleon excited state $N^{*}$ and related physical quantities. Therefore, additional experimental measurements of the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction at threshold are strongly encouraged, and correlation studies could be pursued at facilities such as J-PARC, AMBER, and upcoming HIKE and HIAF meson beam experiments.
I INTRODUCTION
Over the past several decades, meson-nucleon scattering experiments have played a crucial role in the development of hadron spectroscopy wl:2024ext . These experiments are particularly significant, as more than $80\%$ of the light hadrons cataloged by the Particle Data Group (PDG) ParticleDataGroup:2022pth have been discovered or characterized through such studies. Hadron spectroscopy provides an effective means to enhance our understanding of the nonperturbative aspects of strong interactions. In recent years, theorists have also focused on scattering processes involving $K$ and $\pi$ mesons, aiming to explore the production of new hadronic states Liu:2019zoy ; Chen:2016qju ; Chen:2022asf ; Zou:2016bxw ; Lin:2022eau ; Wang:2023lia ; Wang:2022sib ; Xiang:2020phx ; Wang:2019zaw ; Wang:2019uwk ; Wang:2019dsi ; Wang:2018mjz ; Wang:2017qcw ; Cheng:2016ddp ; Wang:2015xwa ; Liu:2021ojf ; Gao:2012zh ; Wu:2009tu ; Xie:2014kja ; Xie:2007qt . Both experimental and theoretical studies suggest that meson beam experiments offer unique advantages in advancing research on hadron spectroscopy and related physical phenomena.
Light-flavor baryons have garnered significant attention from the hadron physics community due to the well-known “missing” problem associated with their excited states ZOU:2005 . Among the numerous observed light-flavor baryon states, the $N^{*}(2080)$ stands out as a particularly intriguing research target, closely linked to the investigation of hidden strange pentaquark states. The nucleon excited state $N^{*}(2080)$ has been proposed as a $K^{*}\Sigma$ molecular state, potentially serving as the strange partner of the $P_{c}^{+}(4457)$ hadronic molecular state He:2017aps ; Lin:2018kcc ; Khemchandani:2011et ; Doring:2008sv . The potential existence of such hadronic molecules has also been suggested in studies utilizing the quark delocalization color screening model Gao:2017hya . As discussed in Ref. wl:2024ext , hidden strange pentaquark states could be identified and studied by measuring processes such as $\pi^{-}p\rightarrow K^{(*)}\Sigma^{(*)}$ and $\pi^{-}p\rightarrow K^{(*)}\Lambda^{(*)}$. These measurements could provide crucial experimental data to address the “missing” problem of baryon excited states. In Ref. Lin:2018kcc , the effective Lagrangian method was employed to calculate the decay modes of the $N^{*}(2080)$ having $J^{P}=3/2^{-}$ as an $S$-wave $K^{*}\Sigma$ molecular state. The results showed that the measured decay properties of $N^{*}(2080)$ with $J^{P}=3/2^{-}$ were well reproduced, supporting the molecular interpretation of this state.
Given that the $J^{P}$ quantum numbers of the $N^{*}(2080)$ [^1] According to the latest PDG ParticleDataGroup:2022pth , the previously identified two-star nucleon resonance $N^{*}(2080)$ has been split into a three-star $N(1875)$ and a two-star $N(2120)$, both with spin-parity $3/2^{-}$. A key characteristic of hadronic molecules is that they are typically very shallowly bound Guo:2017jvc , meaning that their masses are generally just below the corresponding threshold. In this study, we consider the $K^{*}\Sigma$ molecular interpretation, with a threshold of 2086 MeV, and accordingly, we adopt the mass of the $N(2120)$ as 2080 MeV, referring to it by its previous designation, the $N^{*}(2080)$, throughout the remainder of this paper. could be $1/2^{-}$ or $3/2^{-}$, this state can couple to $K^{*}\Sigma$ through $S$ wave interaction. The coupling constant between $N^{*}$ and $K^{*}\Sigma$ can be determined using the Weinberg compositeness criterion Lin:2018kcc ; Weinberg:1965zz ; Baru:2003qq ; Lin:2017mtz . Therefore, searching for and measuring the $N^{*}(2080)$ state via the $\pi^{-}p\rightarrow K^{*}\Sigma$ process is a feasible approach. In the 1970s and 1980s, several experiments measured scattering cross section data for the $\pi^{-}p\rightarrow K^{*}\Sigma$ process Crennell:1972km ; CERN-CollegedeFrance-Madrid-Stockholm:1980mch ; CERN-CollegedeFrance-Madrid-Stockholm:1980ysu ; Dahl:1967pg ; Miller:1965 ; Abramovich:1972rq , providing crucial support for our research.
In this work, we will perform calculations and analysis of the $\pi^{-}p\rightarrow K^{*}\Sigma$ scattering process using the effective Lagrangian approach, with the aim of investigating the production of the nucleon excited state $N^{*}(2080)$ and exploring the constituent counting rule. Additionally, we will analyze the $\pi^{-}p\rightarrow K^{*}\Sigma\rightarrow K^{+}\pi^{-}\Sigma^{0}$ Dalitz process and assess its experimental feasibility, thereby offering valuable theoretical insights for future experimental studies.
Actually, the $\pi^{-}p\rightarrow K^{*}\Sigma$ process serves as an ideal case for examining the constituent field scaling law, commonly known as the constituent counting rule. Initially proposed by Brodsky in 1973 and further developed in the decade following the advent of perturbative QCD (pQCD), this rule embodies the principles of conformality and scale invariance in QCD at high energies Lepage:1980fj ; Mueller:1981sg ; Lepage:1979zb ; Brodsky:1974vy ; Brodsky:1973kr ; Matveev:1973ra ; Guo:2016fqg . The constituent counting rule has been nonperturbatively derived in contexts like AdS/QCD, showcasing its broad applicability to various field theories Polchinski:2001tt . In essence, the rule predicts that in exclusive processes such as $a+b\rightarrow c+d$ involving large momentum transfers, quarks must share significant momenta to coalesce into hadrons by exchanging hard gluons. Consequently, the cross-section for such an exclusive reaction scales according to $d\sigma/dt\sim s^{2-n}f(\theta_{\mathrm{c.m.}})$, where $n=n_{a}+n_{b}+n_{c}+n_{d}$ represents the total number of constituents in particles $a$, $b$, $c$, and $d$. Here, $s$ and $t$ are the Mandelstam variables, and $\theta_{\mathrm{c.m.}}$ is the scattering angle in the center-of-mass frame. The constituent counting rule has been validated by two-body hadronic exclusive reactions conducted at the Brookhaven National Laboratory’s Alternating Gradient Synchrotron (BNL-AGS) Baller:1988tj . More broadly, this rule effectively describes the differential cross-section behavior in various hadron-hadron scattering processes within the large momentum transfer regime JeffersonLabHallA:2002vjd ; JeffersonLabHallA:2004gxp ; White:1994tj ; Amaryan:2021cnj . Applying the constituent counting rule in our analysis of the $\pi^{-}p\rightarrow K^{*}\Sigma$ process not only deepens our understanding of the mechanisms underlying hadron-hadron scattering and the role of constituent quarks but also allows us to assess the accuracy of our theoretical cross-section predictions.
This paper is organized as follows. After the introduction, the Lagrangians and amplitudes used in this work are presented in Section II. The numerical results for the total and differential cross sections are provided in Section III. In Section IV, we introduce the constituent counting rule and assess its validity. The Dalitz process and its experimental feasibility are discussed in Section V, followed by a brief summary in Section VI.
II THE PRODUCTION OF THE $N^{*}(2080)$ VIA THE $\pi^{-}p\rightarrow K^{*}\Sigma$ REACTION
The schematic tree-level Feynman diagrams for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction are depicted in Fig. 1. These include $s$-channel $N^{*}(2080)$ ($\equiv N^{*}$) exchange, $u$-channel $\Sigma$ exchange, and $t$-channel $K$ and $K^{*}$ exchanges.
Figure 1: Feynman diagrams for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction.
To evaluate the pion-induced production of $K^{*}\Sigma$, the relevant Lagrangians are required, as used in previous studies Xiang:2020phx ; Kim:2015ita ; He:2017aps ; Xie:2014kja ; Ben:2023uev ; Wang:2019dsi ; Cheng:2016ddp . In general, similar to the case of the $P_{c}$ pentaquark state, the $J^{P}$ quantum numbers of the $N^{*}(2080)$ state can be either $\frac{3}{2}^{-}$ or $\frac{1}{2}^{-}$. Therefore, for the $s$-channel, both quantum number assignments for the $N^{*}(2080)$ should be considered in our calculations. These Lagrangians include
$\mathcal{L}^{3/2^{-}}_{K^{*}\Sigma N^{*}}$ $=$ $g_{K^{*}\Sigma N^{*}}^{3/2^{-}}\bar{N}^{*}_{\mu}{\Sigma}K^{*\mu}+\mathrm{H.c.},$ (1)
$\mathcal{L}^{3/2^{-}}_{\pi NN^{*}}$ $=$ $\frac{g_{\pi NN^{*}}^{3/2^{-}}}{m_{\pi}}\bar{N}\gamma_{5}\bm{\tau}\cdot\partial^{\mu}\bm{\pi}N^{*}_{\mu}+\mathrm{H.c.},$ (2)
$\mathcal{L}^{1/2^{-}}_{K^{*}\Sigma N^{*}}$ $=$ $g_{K^{*}\Sigma N^{*}}^{1/2^{-}}\bar{N}^{*}\gamma_{5}\gamma^{\mu}K^{*}_{\mu}\Sigma+\mathrm{H.c.},$ (3)
$\mathcal{L}^{1/2^{-}}_{\pi NN^{*}}$ $=$ $g_{\pi NN^{*}}^{1/2^{-}}\bar{N}\bm{\tau}\cdot\bm{\pi}N^{*}+\mathrm{H.c.}.$ (4)
The Lagrangians corresponding to the $u$-channel are written as
$\mathcal{L}_{\Sigma\Sigma\pi}$ $=$ $-\frac{g_{\Sigma\Sigma\pi}}{m_{\pi}}\bar{\Sigma}\gamma^{5}\gamma^{\mu}\partial_{\mu}\bm{\pi}\cdot\bm{T}\Sigma,$ (5)
$\mathcal{L}_{K^{*}N{\Sigma}}$ $=$ $-g_{K^{*}N{\Sigma}}\bar{N}\Sigma(K\mkern-11.0mu/^{*}-\frac{k_{K^{*}N{\Sigma}}}{2m_{N}}\sigma_{\mu\nu}\partial^{\nu}K^{*\mu})+\mathrm{H.c.},$ (6)
The effective Lagrangians for the $t$-channel involving $K$ and $K^{*}$ exchanges are given as follows:
$\mathcal{L}_{\pi KK^{*}}$ $=$ $g_{\pi KK^{*}}[\bar{K}(\partial^{\mu}\bm{\tau}\cdot\bm{\pi})-(\partial^{\mu}\bar{K})\bm{\tau}\cdot\bm{\pi}]K^{*}_{\mu}+\mathrm{H.c.},$ (7)
$\mathcal{L}_{\pi K^{*}K^{*}}$ $=$ $g_{\pi K^{*}K^{*}}\epsilon^{\mu\nu\alpha\beta}\partial_{\mu}\bar{K}^{*}_{\nu}\bm{\tau}\cdot\bm{\pi}\partial_{\alpha}K^{*}_{\beta},$ (8)
$\mathcal{L}_{KN\Sigma}$ $=$ $g_{KN\Sigma}\bar{N}\gamma_{5}K\bm{\tau}\cdot\bm{\Sigma}+\mathrm{H.c.},$ (9)
$\mathcal{L}_{K^{*}N{\Sigma}}$ $=$ $-g_{K^{*}N{\Sigma}}\bar{N}\Sigma(K\mkern-11.0mu/^{*}-\frac{k_{K^{*}N{\Sigma}}}{2m_{N}}\sigma_{\mu\nu}\partial^{\nu}K^{*\mu})+\mathrm{H.c.},$ (10)
where $\bm{\tau}$ denotes the Pauli matrices, and the matrix $\bm{T}$ is defined as in Ref. Matsuyama:2006rp . The symbols $\bm{\pi}$, $K$, and $K^{*}$ represent the $\pi$, $K$, and $K^{*}$ meson fields, respectively. Similarly, $N$, $\Sigma$, and $N^{*}$ correspond to the nucleon, $\Sigma$ baryon, and $N^{*}(2080)$ fields, respectively.
Assuming the $N^{*}(2080)(\frac{3}{2}^{-})$ is a pure $S$-wave molecular state of $K^{*}$ and $\Sigma$, the coupling constant $g_{K^{*}\Sigma N^{*}}^{3/2^{-}}$ can be estimated model-independently using the Weinberg compositeness criterion, yielding Weinberg:1965zz ; Baru:2003qq ; Lin:2017mtz ; Lin:2018kcc
$g_{K^{*}\Sigma N^{*}}^{2}$ $=$ $\frac{4\pi}{4M_{N^{*}}M_{\Sigma}}\frac{(M_{K^{*}}+M_{\Sigma})^{5/2}}{(M_{K^{*}}M_{\Sigma})^{1/2}}\sqrt{32\epsilon},$ (11)
$\epsilon$$\equiv$ $M_{K^{*}}+M_{\Sigma}-M_{N^{*}},$ (12)
where $M_{N^{*}}$, $M_{K^{*}}$, and $M_{\Sigma}$ represent the masses of the $N^{*}(2080)(\frac{3}{2}^{-})$, $K^{*}$, and $\Sigma$, respectively, while $\epsilon$ denotes the $K^{*}\Sigma$ binding energy. Assuming the physical state is a pure $S$-wave hadronic molecule, the relative uncertainty of the above approximation for the coupling constant is given by $\sqrt{2\mu\epsilon}\cdot r$, where $\mu=\frac{m_{1}m_{2}}{m_{1}+m_{2}}$ is the reduced mass of the bound particles, and $r$ is the range of forces, which can be estimated as the inverse of the mass of the exchanged particle. For the $K^{*}\Sigma$ system, $r$ can be estimated as $1/m_{K}$ Lin:2018kcc .
Following Ref. Lin:2018kcc , we adopt the mass of the $N^{*}(2080)(\frac{3}{2}^{-})$ as $M_{N^{*}}=2080\text{ MeV}$. Substituting this into Eq. (11), one obtains:
$g_{K^{*}\Sigma N^{*}}^{3/2^{-}}=1.72.$ (13)
Since the $N^{*}(2080)(\frac{1}{2}^{-})$ can also be considered a $S$-wave molecular state of $K^{*}$ and $\Sigma$, we approximate the value of $g_{K^{*}\Sigma N^{*}}^{1/2^{-}}$ as 1.72. In the absence of experimental measurements for the partial width of the $N^{*}(2080)$ decays to $\pi N$, we treat the coupling constants $g_{\pi NN^{*}}^{3/2^{-}}$ and $g_{\pi NN^{*}}^{1/2^{-}}$ as free parameters to be determined through fitting.
The coupling constant values for the $\Sigma\Sigma\pi$, $K^{*}N\Sigma$, and $KN\Sigma$ interactions have been provided in several theoretical works Xiang:2020phx ; He:2017aps ; Cheng:2016ddp . We use the following values: $g_{\Sigma\Sigma\pi}=0.8$, $g_{K^{*}N\Sigma}=-2.46$, $\kappa_{K^{*}N\Sigma}=-0.47$, and $g_{KN\Sigma}=2.69$.
The coupling constant $g_{\pi KK^{*}}$ can be determined from the decay width $\Gamma_{K^{*}\to K\pi}$, as detailed in Xiang:2020phx
${\Gamma}_{K^{*}{\rightarrow}K\pi}$ $=$ $\frac{g^{2}_{\pi KK^{*}}}{2\pi}\frac{|\vec{p}_{\pi}^{c.m.}|^{3}}{m^{2}_{K^{*}}}$ (14)
with
$|\vec{p}_{\pi}^{c.m.}|$ $=$ $\frac{\sqrt{[m^{2}_{K^{*}}-(m_{K}+m_{\pi})^{2}][m^{2}_{K^{*}}-(m_{K}-m_{\pi})^{2}]}}{2m_{K^{*}}},$ (15)
[…]
For the $t$-channel and $u$-channel, the form factors are adopted in our calculation as follows,
$F_{t/u}(q)=\frac{\Lambda{{}^{2}_{t/u}}-m^{2}}{\Lambda{{}^{2}_{t/u}}-q^{2}}.$ (16)
To account for the finite size of the hadron, we incorporate a form factor in our calculation for the $s$-channel with an intermediate baryon, as follows:
$F_{s}(q)=\frac{\Lambda{{}^{4}_{s}}}{\Lambda{{}^{4}_{s}}+(q^{2}-m^{2})^{2}},$ (17)
where $q$ and $m$ represent the four-momentum and mass of the exchanged particles, respectively.
Based on the above Lagrangians, the scattering amplitude for the reaction $\pi^{-}p\rightarrow K^{*}\Sigma$ is given by:
$\mathcal{M}$ $=$ $\epsilon^{\mu}(k_{2})\bar{u}(p_{2})(\mathcal{A}_{s,\mu}+\mathcal{A}_{u,\mu}+\mathcal{A}_{t,\mu}){u}(p_{1}),$ (18)
where $\epsilon^{\mu}$ denotes the polarization vector of the outgoing $K^{*}$ meson, while $\bar{u}$ and $u$ represent the Dirac spinors for the outgoing $\Sigma$ baryon and the incoming proton, respectively.
The reduced amplitudes $\mathcal{A}_{s,\mu}$, $\mathcal{A}_{u,\mu}$, and $\mathcal{A}_{t,\mu}$ for the $s$-, $u$-, and $t$-channel contributions read as
$\mathcal{A}^{N^{*({3/2}^{-})}}_{s,\mu}$ $=$ $-i\sqrt{2}g_{K^{*}\Sigma{N^{*}}}^{{3/2}^{-}}\frac{g_{\pi N{N^{*}}}^{{3/2}^{-}}}{m_{\pi}}F_{s}(q)\gamma_{5}k_{1}^{\nu}$ (19)
×$\times\frac{(q\mkern-8.0mu/_{s}+m_{N^{*}})}{s-m^{2}_{N^{*}}+im_{N^{*}}\Gamma_{N^{*}}}\Delta_{\mu\nu},$
$\mathcal{A}^{N^{*({1/2}^{-})}}_{s,\mu}$ $=$ $-\sqrt{2}g_{K^{*}\Sigma{N^{*}}}^{{1/2}^{-}}g_{\pi N{N^{*}}}^{{1/2}^{-}}F_{s}(q)\gamma_{5}\gamma_{\mu}$ (20)
×$\times\frac{(q\mkern-8.0mu/_{s}+m_{N^{*}})}{s-m^{2}_{N^{*}}+im_{N^{*}}\Gamma_{N^{*}}},$
$\mathcal{A}^{\Sigma}_{u,\mu}$ $=$ $-\sqrt{2}\frac{g_{\Sigma\Sigma\pi}}{m_{\pi}}g_{K^{*}N\Sigma}F_{u}(q)\gamma^{5}\gamma^{\nu}k_{1\nu}\frac{(q\mkern-8.0mu/_{\Sigma}+m_{\Sigma})}{{u-m_{\Sigma}^{2}}}$ (21)
×$\times(\gamma_{\mu}-\frac{k_{K^{*}{\Sigma}N}}{2m_{N}}\gamma_{\mu}q\mkern-8.0mu/_{K^{*}}),$
$\mathcal{A}^{K}_{t,\mu}$ $=$ $-\sqrt{2}g_{K^{*}K\pi}g_{K\Sigma N}F_{t}(q)\frac{1}{{t-m_{K}^{2}}}\gamma_{5}$ (22)
×$\times[k_{1\mu}+(k_{1}-k_{2})_{\mu}],$
$\mathcal{A}^{K^{*}}_{t,\mu}$ $=$ $-i\sqrt{2}g_{\pi K^{*}K^{*}}g_{K^{*}N\Sigma}F_{t}(q)(\gamma_{\xi}-\frac{k_{K^{*}{\Sigma}N}}{2m_{N}}\gamma_{\xi}q\mkern-8.0mu/_{K^{*}})$ (23)
×$\times\epsilon_{\mu\nu\alpha\beta}\frac{\mathcal{P}^{\nu\xi}}{{t-m_{K^{*}}^{2}}}k_{2}^{\alpha}(k_{1}-k_{2})^{\beta}$
with
$\mathcal{P}^{\nu\xi}$ $=$ $i(g^{\nu\xi}+q_{K^{*}}^{\nu}q_{K^{*}}^{\xi}/{m_{K^{*}}^{2}}),$ (24)
$\Delta_{\mu\nu}$ $=$ $-g_{\mu\nu}+\frac{1}{3}\gamma_{\mu}\gamma_{\nu}$ (25)
$+\frac{1}{3m_{N^{*}}}(\gamma_{\mu}q_{\nu}-\gamma_{\nu}q_{\mu})+\frac{2}{3m_{N^{*}}^{2}}q_{\mu}q_{\nu},$
where $s=(k_{1}+p_{1})^{2}$, $u=(p_{2}-k_{1})^{2}$, and $t=(k_{1}-k_{2})^{2}$ are the Mandelstam variables.
The Regge trajectory model has been successful in analyzing hadron production at intermediate and high energies Wang:2015hfm ; Ozaki:2009wp ; Wang:2015zcp ; Wang:2017qcw ; Storrow:1983ct . In this model, Reggeization is achieved by replacing the $t$-channel propagator in the Feynman amplitudes (Eqs. (22) and (23)) with the Regge propagator
$\frac{1}{t-m_{K}^{2}}\rightarrow(\frac{s}{s_{scale}})^{\alpha_{K}(t)}\frac{\pi\alpha^{\prime}_{K}}{\Gamma[1+{\alpha_{K}(t)}]\sin[\pi{\alpha_{K}(t)}]},$ (26)
$\frac{1}{t-m_{K^{*}}^{2}}\rightarrow(\frac{s}{s_{scale}})^{\alpha_{K^{*}}(t)-1}\frac{\pi\alpha^{\prime}_{K^{*}}}{\Gamma[{\alpha_{K^{*}}(t)}]\sin[\pi{\alpha_{K^{*}}(t)}]}.$ (27)
The scale factor $s_{\text{scale}}$ is fixed at 1 GeV. The Regge trajectories for $\alpha_{K}(t)$ and $\alpha_{K^{*}}(t)$ are given as Ozaki:2009wp :
$\alpha_{K}(t)=0.70(t-m_{K}^{2}),\alpha_{K^{*}}(t)=1+0.85(t-m_{K^{*}}^{2}).$ (28)
Notably, no additional free parameters are introduced with the incorporation of the Regge model.
III Numerical results
After completing the preparation, the differential cross section for the reaction $\pi^{-}p\rightarrow K^{*}\Sigma$ can be calculated and used for correlation analysis with experimental data. In the center of mass (c.m.) frame, the differential cross section is given by:
$\frac{d\sigma}{d\cos\theta}=\frac{1}{32\pi s}\frac{|\vec{k}_{2}^{c.m.}|}{|\vec{k}_{1}^{c.m.}|}\left(\frac{1}{2}\sum\limits_{\lambda}|\mathcal{M}|^{2}\right),$ (29)
where $\theta$ is the angle between the outgoing $K^{*}$ meson and the direction of the $\pi$ beam in the c.m. frame. $\vec{k}_{1}^{c.m.}$ and $\vec{k}_{2}^{c.m.}$ denote the three-momenta of the initial $\pi$ beam and the final $K^{*}$ meson, respectively.
III.1 Production of the $N^{*}(2080)$ with $J^{P}$ = $\frac{3}{2}^{-}$
Using the MINUIT code from CERNLIB, we will fit the experimental data CERN-CollegedeFrance-Madrid-Stockholm:1980ysu ; CERN-CollegedeFrance-Madrid-Stockholm:1980mch ; Crennell:1972km ; Abramovich:1972rq ; Dahl:1967pg ; Miller:1965 for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction. Both total and differential cross section datas are used in a $\chi^{2}$ fitting algorithm to determine the values of the free parameters. The fitting parameters are listed in Table 1. The fit excluding the contribution of the $N^{*}(2080)(\frac{3}{2}^{-})$ yields $\chi^{2}/d.o.f.$ = 6.19, while including the $N^{*}(2080)(\frac{3}{2}^{-})$ improves the fit to $\chi^{2}/d.o.f.$ = 1.52. This indicates a significant contribution from the $s$-channel $N^{*}(2080)(\frac{3}{2}^{-})$.
Table 1: Fitted values of the free parameters with all experimental data in Ref. CERN-CollegedeFrance-Madrid-Stockholm:1980ysu ; CERN-CollegedeFrance-Madrid-Stockholm:1980mch ; Crennell:1972km ; Abramovich:1972rq ; Dahl:1967pg ; Miller:1965 for the case of the $N^{*}(2080)$ with $J^{P}=\frac{3}{2}^{-}$.
with $N^{*}(2080)(\frac{3}{2}^{-})$ without $N^{*}(2080)(\frac{3}{2}^{-})$
$\Lambda_{K}$ 1.53$\pm$0.04 1.90$\pm$0.01
$\Lambda_{K^{*}}$ 1.63$\pm$0.05 1.85$\pm$0.07
$\Lambda_{u}$ 1.19$\pm$0.05 1.13$\pm$0.01
$\Lambda_{s}$ 1.18$\pm$0.06
$g_{\pi N{N^{*}}}^{{3/2}^{-}}$ 0.37$\pm$0.04
$\chi^{2}/d.o.f.$ 1.52 6.19
The total cross section for the reaction $\pi^{-}p\rightarrow K^{*}\Sigma$ with the $N^{*}(2080)$ having $J^{P}$ = $\frac{3}{2}^{-}$ is illustrated in Fig. 2. The results, which include the contributions from the $t$-channel, $u$-channel, and $s$-channel, show good agreement with the experimental data. Additionally, Fig. 2 demonstrates that while the contributions from the $t$- and $u$-channels are relatively small, the $s$-channel plays a crucial role in the reaction.
Figure 2: The total cross section for the reaction $\pi^{-}p\rightarrow K^{*}\Sigma$. The band stands for the error bar of the five fitting parameters in Table 1. The solid (red), dashed-dotted (blue), dashed (purple), dash–double dotted (green) lines are for the full model, the $s$-channel, the $t$-channel and the $u$-channel, respectively. Here, the spin-parity quantum number of the $N^{*}(2080)$ is $\frac{3}{2}^{-}$.
Figure 3: The differential cross-section $d\sigma/dcos\theta$ for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction as a function of $\cos\theta$. The experimental datas are from Ref. Crennell:1972km . Here, the notation is the same as that in Fig. 2.
In Fig. 3, we present the differential cross section for the reaction $\pi^{-}p\rightarrow K^{*}\Sigma$ as a function of $\cos\theta$, specifically for the $N^{*}(2080)$ with $J^{P}=\frac{3}{2}^{-}$ at a center-of-mass energy $W=3.056$ GeV. The analysis reveals that the $t$-channel contribution dominates at forward angles, while the $u$-channel contribution is small and negligible. In Fig. 4, we show the differential cross section for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction as a function of center-of-mass energy, with $\cos\theta=1$, for the same $N^{*}(2080)$ with $J^{P}=\frac{3}{2}^{-}$. Since these measurements are taken at forward angles, the distinct shapes of the curves resulting from the contributions of the three channels allow us to effectively isolate the contributions from the $t$-channel and $u$-channel, while providing a clear indication of the $s$-channel $N^{*}(2080)(\frac{3}{2}^{-})$. This approach is ideal for studying the contribution of the $N^{*}(2080)(\frac{3}{2}^{-})$ through differential cross sections at forward angles.
In Fig. 5, we calculate the differential cross section of the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction at various center-of-mass energies. The results indicate that as the energy approaches the threshold of $K^{*}\Sigma$, the contribution of the $s$-channel becomes more prominent. Conversely, as energy increases, the contribution of the $t$-channel at forward angles becomes more significant, while the $u$-channel contribution remains consistently small. These findings suggest that measuring the differential cross section of the reaction $\pi^{-}p\rightarrow K^{*}\Sigma$ near the $K^{*}\Sigma$ threshold will provide valuable information about the $N^{*}(2080)(\frac{3}{2}^{-})$ resonance.
Figure 4: The differential cross-section $d\sigma/d\cos\theta$ for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction varies with different c.m. energies when $\cos\theta$ = 1. Here, the notation is the same as that in Fig. 2.
Figure 5: The differential cross section $d\sigma/d\cos\theta$ of the $\pi^{-}p\rightarrow K^{*}\Sigma$ process as a function of $\cos\theta$ at different c.m. energies. Here, the notation is the same as that in Fig. 2.
The $t$-distribution for the reaction $\pi^{-}p\rightarrow K^{*}\Sigma$ with the $N^{*}(2080)$ and $J^{P}=\frac{3}{2}^{-}$ at $W=2.88$ GeV is shown in Fig. 6. The results, incorporating full contributions from the $s$-channel, $t$-channel, and $u$-channel, are in good agreement with the experimental data. Fig. 7 presents the $t$-distribution for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction at various center-of-mass energies. These results provide insight into the production mechanism of the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction, aiding in the differentiation of its underlying processes.
Figure 6: The $t$ distribution for the reaction $\pi^{-}p\rightarrow K^{*}\Sigma$. The experimental datas are from Ref. CERN-CollegedeFrance-Madrid-Stockholm:1980mch . Here, the notation is the same as that in Fig. 2.
Figure 7: The $t$-distribution for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction at different c.m. energies $W$ = 2.2 GeV, 2.3 GeV, 2.4 GeV and 2.5 GeV. Here, the notation is the same as in Fig. 2.
III.2 Production of the $N^{*}(2080)$ with $J^{P}$ = $\frac{1}{2}^{-}$
Similar to the analysis of $N^{*}(2080)$ with $J^{P}$ = $\frac{3}{2}^{-}$, we also calculate the cross section for the reaction $\pi^{-}p\rightarrow K^{*}\Sigma$ with $N^{*}(2080)$ having $J^{P}$ = $\frac{1}{2}^{-}$ as detailed in Figs. 8, 9 and 10. We employ the same fitting scheme as in Table 1, treating the cutoff value as a free parameter. The fitting parameters are shown in Table 2. The fit excluding the $N^{*}(2080)(\frac{1}{2}^{-})$ contribution yields a $\chi^{2}/d.o.f.$ = 6.19, whereas including this contribution results in $\chi^{2}/d.o.f.$ = 2.20. For the $N^{*}(2080)$ with $J^{P}$ = $\frac{3}{2}^{-}$, the $\chi^{2}/d.o.f.$ = 1.52, indicating a better fit to the experimental data for the $N^{*}(2080)(\frac{3}{2}^{-})$.
Fig. 9 and Fig. 10 present the differential cross section $d\sigma/d\cos\theta$ and the $t$-distribution for $\pi^{-}p\rightarrow K^{*}\Sigma$, respectively. These results closely match those in Fig. 3 and Fig. 6. Fig. 8 shows the total cross section for $\pi^{-}p\rightarrow K^{*}\Sigma$, with a peak value approximately 500 $\mu b$ higher than in Fig. 2. In Fig. 11, the differential cross section for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction as a function of c.m. energies at $\cos\theta$ = 1 is presented. This provides an ideal means to assess the contribution of the $N^{*}(2080)(\frac{1}{2}^{-})$ through forward-angle differential cross sections. Fig. 12 shows the differential cross section at various c.m. energies. As the energy approaches the $K^{*}\Sigma$ threshold, the $s$-channel contribution becomes more pronounced. With increasing energy, the $t$-channel contribution at forward angles grows more significant, while the $u$-channel contribution remains small. This behavior is similar to that observed for the $N^{*}(2080)(\frac{3}{2}^{-})$. Fig. 13 displays the $t$-distribution for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction at different c.m. energies. These results may aid us in distinguishing the production mechanisms for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction with the $N^{*}(2080)$ having $J^{P}$ = $\frac{1}{2}^{-}$.
Table 2: Fitted values of the free parameters with all experimental data in Refs. CERN-CollegedeFrance-Madrid-Stockholm:1980ysu ; CERN-CollegedeFrance-Madrid-Stockholm:1980mch ; Crennell:1972km ; Abramovich:1972rq ; Dahl:1967pg ; Miller:1965 for the case of the $N^{*}(2080)$ with $J^{P}=\frac{1}{2}^{-}$.
with $N^{*}(2080)(\frac{1}{2}^{-})$ without $N^{*}(2080)(\frac{1}{2}^{-})$
$\Lambda_{K}$ 1.81 $\pm$0.02 1.90$\pm$0.01
$\Lambda_{K^{*}}$ 1.71$\pm$0.02 1.85$\pm$0.07
$\Lambda_{u}$ 1.14$\pm$0.01 1.13$\pm$0.01
$\Lambda_{s}$ 1.61$\pm$0.02
$g_{\pi N{N^{*}}}^{1/2^{-}}$ 0.20$\pm$0.01
$\chi^{2}/d.o.f.$ 2.20 6.19
Figure 8: The total cross section for the reaction $\pi^{-}p\rightarrow K^{*}\Sigma$. The band stands for the error bar of the five fitting parameters in Table 2. The solid (purple), dashed-dotted (blue), dashed (red), dash–double dotted (green) lines are for the full model, the $s$-channel, the $t$-channel and the $u$-channel, respectively. Here, the spin-parity quantum number of the $N^{*}(2080)$ is $\frac{1}{2}^{-}$.
Figure 9: The differential cross-section $d\sigma/d\cos\theta$ for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction as a function of $\cos\theta$. The experimental datas are from Ref. Crennell:1972km . Here, the notation is the same as that in Fig. 8.
Figure 10: The $t$ distribution for the reaction $\pi^{-}p\rightarrow K^{*}\Sigma$. The experimental datas are from Ref. CERN-CollegedeFrance-Madrid-Stockholm:1980mch . Here, the notation is the same as that in Fig. 8.
Figure 11: The differential cross-section $d\sigma/d\cos\theta$ for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction varies with different c.m. energies when $\cos\theta$ = 1. Here, the notation is the same as that in Fig. 8.
Figure 12: The differential cross section $d\sigma/d\cos\theta$ of the $\pi^{-}p\rightarrow K^{*}\Sigma$ process as a function of $\cos\theta$ at different c.m. energies. Here, the notation is the same as that in Fig. 8.
Figure 13: The $t$-distribution for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction at different c.m. energies $W$ = 2.2 GeV, 2.3 GeV, 2.4 GeV and 2.5 GeV. Here, the notation is the same as in Fig. 8.
IV constituent counting rule in hard exclusive reactions
We introduce the constituent counting rule for hard exclusive reactions by presenting the expressions for the cross section and matrix element, and by explaining a typical hard gluon exchange process that contributes to such reactions. For a large-angle exclusive scattering process $a+b\rightarrow c+d$, the reaction cross section is given by:
$\frac{d\sigma_{ab\to cd}}{dt}\simeq\frac{1}{16\pi s^{2}}\overline{\sum_{\mathrm{pol}}}|M_{ab\to cd}|^{2},$ (30)
In this context, the approximation implies that particle masses are neglected. The Mandelstam variables $s$ and $t$ are defined as follows:
[…]
where $p_{i}$ ($i=a,b,c,d$) represents the momentum of hadron $i$, while $p_{\mathrm{c.m.}}$ and $\theta_{\mathrm{c.m.}}$ denote the momentum and scattering angle in the center of mass (c.m.) frame, respectively. Eq. (30) includes an average over the initial spins and a sum over the final spins.
The matrix element $M_{ab\rightarrow cd}$ can be expressed in a factorized form when there is a large momentum transfer Lepage:1980fj ; Mueller:1981sg ; Lepage:1979zb ; Efremov:1979qk ; Guo:2016fqg :
[…]
Here, $[x]$ represents the set of light-cone momentum fractions of partons within a hadron, where $x_{i}=p^{+}_{i}/p^{+}$, with $p_{i}$ and $p$ denoting the momenta of the $i$th parton and the hadron, respectively. The light-cone component is defined as $p^{+}=(p^{0}+p^{3})/\sqrt{2}$, considering the third axis as the longitudinal direction.
For a nucleon, two independent variables are required to describe the distribution amplitude, while accounting for the constraint of momentum conservation, $x_{1}+x_{2}+x_{3}=1$. Thus, we have $[x]=x_{1},x_{2}$ and $[dx]=dx_{1}dx_{2}$ in Eq. (32). Conversely, only one variable, $x$, is needed for mesons such as pions and kaons. The variable $Q^{2}$ represents the hard scale of the reaction, approximately given by $Q^{2}\simeq s$ for large-angle elastic exclusive scattering. In Eq. (32), the dependencies on the renormalization and factorization scales are suppressed: the former is governed by the renormalization group equation for the coupling constant, and the latter by the Efremov-Radyushkin–Brodsky–Lepage (ERBL) evolution equations for the distribution amplitudes Lepage:1979zb ; Efremov:1979qk . These scales are typically taken to be on the order of $Q^{2}$ to minimize large radiative corrections. Consequently, the $Q^{2}$ dependencies are logarithmic and do not significantly affect the scaling behavior of the matrix element.
Originally, the constituent counting rule was proposed based on dimensional counting Brodsky:1973kr ; Brodsky:1974vy ; Matveev:1973ra . It was subsequently explored within the framework of perturbative QCD, incorporating hard scattering processes Farrar:1979aw ; Efremov:1978rn ; Duncan:1979hi ; Duncan:1980qd ; Lepage:1980fj . In this section, we derive the scaling rule by counting mass dimensions.
Given that the state vector of a hadron is normalized as $\langle h(p^{\prime})|h(p)\rangle=2p^{0}(2\pi)^{3}\delta^{(3)}(\vec{p}-\vec{p}^{\prime})$, we find that its mass dimension is $[1/M]$. If a hadron is composed of $n_{l}$ elementary constituents, its state vector can be expressed as
$|l\rangle=\sqrt{N_{l}}|n_{l}\rangle,\quad\left[\sqrt{N_{l}}\right]=\left[M^{n_{l}-1}\right].$ (33)
The second equation indicates that the normalization factor $\sqrt{N_{l}}$ carries the mass dimension $\left[M^{n_{l}-1}\right]$ when the state vector of each constituent has the mass dimension $[1/M]$. In practice, this normalization factor corresponds to the “decay constant,” which is independent of the hard momentum scale Kawamura:2013iia .
We now consider the mass dimensions of the matrix element in Eq. (30) to derive the counting rule for exclusive cross sections. The scattering matrix $S$ is connected to the transition matrix $T$ by the relation $S=\bm{1}+i(2\pi)^{4}\delta^{(4)}(p_{f}-p_{i})T$, indicating that the mass dimension of $T$ is $[T]=[M^{4}]$. The matrix element $M_{ab\rightarrow cd}$ is then expressed in terms of $T$ as follows:
$M_{ab\to cd}=\langle cd|T|ab\rangle=\sqrt{N_{a}N_{b}N_{c}N_{d}}\langle n_{c}n_{d}|T|n_{a}n_{b}\rangle.$ (34)
Since the normalization factors $N_{i}(i=a,b,c,d)$ are represented by soft constants, we analyze the matrix element $\hat{M}_{ab\rightarrow cd}$ by omitting these factors. Consequently, the remaining hard component should be expressed in terms of the variables $s$ and $t$:
[…]
Given the dimensions $[T]=[M^{4}]$ and $[|n_{i}\rangle]=[1/M^{n_{i}}]$, the dimension of the matrix element is given by
$\begin{bmatrix}\hat{M}_{ab\to cd}\end{bmatrix}=\begin{bmatrix}\langle n_{c}n_{d}|T|n_{a}n_{b}\rangle\end{bmatrix}=\begin{bmatrix}M^{4-n}\end{bmatrix},$ (36)
where $n\equiv n_{a}+n_{b}+n_{c}+n_{d}$. In a large-angle exclusive reaction, the variable $s$ can be chosen as the sole hard scale. Thus, considering the mass dimension, the matrix element is expressed as:
$\hat{M}_{ab\to cd}=\hat{F}_{ab\to cd}(s,t)=s^{(4-n)/2}F_{ab\to cd}(t/s),$ (37)
where $F_{ab\to cd}(t/s)$ is a dimensionless function of scattering angle, with $-2t/s=1-\cos\theta_{c.m.}$ from Eq. (31). Using Eqs. (30), (35), and (37), we derive the constituent-counting expression for the cross section:
$\frac{d\sigma_{ab\to cd}}{dt}=\frac{1}{s^{n-2}}f_{ab\to cd}(t/s),$ (38)
where $f(t/s)$ accounts for the scattering angle and is multiplied by the normalization factors. The mass dimensions of $f(t/s)$ are given by $[f(t/s)]=[N_{a}N_{b}N_{c}N_{d}|F_{ab\to cd}|^{2}]=[M^{2n-8}]$, so the overall mass dimension of Eq. (38) is $[1/M^{4}]$. This derivation illustrates the counting rule based on mass dimensions, referred to as the “constituent counting rule” since it quantifies the number of constituents involved in the reaction.
Experimentally, the constituent counting rule was first applied in Ref. Anderson:1976ph to reactions such as $\gamma p\rightarrow\pi^{+}n$, $\pi^{0}p$, $\pi^{-}\Delta^{++}$, $\rho^{0}p$, $K^{+}\Lambda$, and $K^{+}\Sigma^{0}$, suggesting $n=3$ for baryons and $n=2$ for mesons. The JLab (Thomas Jefferson National Accelerator Facility) Hall-A and E94-104 collaborations subsequently verified that experimental data are consistent with $n_{\gamma}=1$, $n_{p}=n_{n}=3$, and $n_{\pi^{+}}=2$ by analyzing the $\gamma p\rightarrow\pi^{+}n$ reaction JeffersonLabHallA:2002vjd ; JeffersonLabHallA:2004gxp . Furthermore, several experiments have confirmed the scaling behavior of the exclusive cross section predicted by the constituent counting rule Baller:1988tj ; White:1994tj .
Table 3: The cross sections of $\pi^{-}p\rightarrow K^{*}\Sigma$ at $\theta_{\mathrm{c.m.}}=90^{\circ}$ for the case of the $N^{*}(2080)$ with $J^{P}$ = $\frac{3}{2}^{-}$.
$\sqrt{s}$ (GeV) $t$ $(\mathrm{GeV}^{2})$ $\frac{d\sigma}{dt}$ $(\mu b/{\mathrm{GeV}}^{2})$
2.30 1.135 80.655$\pm$24.405
2.35 1.249 55.077$\pm$18.164
2.40 1.366 37.173$\pm$13.104
2.45 1.485 25.063$\pm$9.296
2.50 1.607 17.010$\pm$6.650
2.55 1.732 11.673$\pm$4.640
2.60 1.859 8.121$\pm$3.306
2.65 1.989 5.740$\pm$2.381
2.70 2.121 4.127$\pm$1.738
2.75 2.256 3.021$\pm$1.288
2.80 2.393 2.250$\pm$0.970
Table 4: The cross sections of $\pi^{-}p\rightarrow K^{*}\Sigma$ at $\theta_{\mathrm{c.m.}}=90^{\circ}$ for the case of the $N^{*}(2080)$ with $J^{P}$ = $\frac{1}{2}^{-}$.
$\sqrt{s}$ (GeV) $t$ (GeV2) $\frac{d\sigma}{dt}$ $(\mu b/{\mathrm{GeV}}^{2})$
2.30 1.135 104.376$\pm$14.620
2.35 1.249 71.407$\pm$8.284
2.40 1.366 48.478$\pm$5.898
2.45 1.485 33.337$\pm$4.257
2.50 1.607 23.150$\pm$3.098
2.55 1.732 16.212$\pm$2.267
2.60 1.859 11.444$\pm$1.667
2.65 1.989 8.144$\pm$1.230
2.70 2.121 5.844$\pm$0.912
2.75 2.256 4.230$\pm$0.760
2.80 2.393 3.122$\pm$0.509
In this study, we test the constituent counting rule using the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction. According to the rule, the total number of constituents is $n=2+3+2+3=10$. We calculate the cross sections for the $\pi^{-}p\rightarrow K^{*}\Sigma$ process at $\theta_{\mathrm{c.m.}}=90^{\circ}$, considering the cases where the $N^{*}(2080)$ has $J^{P}=\frac{3}{2}^{-}$ and $J^{P}=\frac{1}{2}^{-}$. The results are presented in Table 3 and Table 4, respectively.
By fitting the experimental data at $\sqrt{s}=2.88$ GeV and the numerical results shown in Table 3 and Table 4 to the expression $d\sigma/dt=(\text{constant})\times s^{2-n}$, the scaling factors obtained are presented in Table 5. We find that for both $N^{*}(2080)(\frac{3}{2}^{-})$ and $N^{*}(2080)(\frac{1}{2}^{-})$, the fitted values of $n$ are very close to 10, confirming the accuracy of the constituent counting rule. Fig. 14 and Fig. 15 depict the differential cross section $d\sigma/dt$ of the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction at a large meson production angle ($\theta=90^{\circ}$ in the center of mass frame) as a function of $\sqrt{s}$ for the $N^{*}(2080)(\frac{3}{2}^{-})$ and $N^{*}(2080)(\frac{1}{2}^{-})$, respectively.
Table 5: Fitted values of the scaling factor $n$ by using the $\chi^{2}$ fitting algorithm.
$N^{*}(2080)(\frac{3}{2}^{-})$ $N^{*}(2080)(\frac{1}{2}^{-})$
$n$ 10.01$\pm$0.01 10.40$\pm$0.04
$\chi^{2}/d.o.f.$ 0.76 0.99
Figure 14: Differential cross section of $\pi^{-}p\rightarrow K^{*}\Sigma$, $d\sigma/dt$, at large meson production angle $\theta=90^{\circ}$ in c.m. as a function of $W$ $(W=|\sqrt{s}|)$. The green triangle and the black squares represent the experimental point and the numerical points in Table 3, respectively. The band stands for the error bar of the fitting parameters in Table 5.
Figure 15: Differential cross section of $\pi^{-}p\rightarrow K^{*}\Sigma$, $d\sigma/dt$, at large meson production angle $\theta=90^{\circ}$ in c.m. as a function of $(W=|\sqrt{s}|)$. The green triangle and the black squares represent the experimental point and the numerical points in Table 4, respectively. The band stands for the error bar of the fitting parameters in Table 5.
In this work, the constituent counting rule of perturbative QCD could be utilized to determine the internal quark configuration of exotic-hadron candidates in exclusive production processesKawamura:2013iia . In the exclusive process $a+b\rightarrow c+d$ with large-momentum transfer, quarks should share large momenta so that they can stick together to form a hadron by exchanging hard gluons. Since the cross section of the $a+b\rightarrow c+d$ exclusive reaction should scale like $d\sigma/dt\sim s^{2-n}f(\theta_{\mathrm{c.m.}})$ with $n=n_{a}+n_{b}+n_{c}+n_{d}$, where $n_{i}(i=a,b,c,d)$ is the number of constituents in the particle $i$. Since the factor $n_{i}$ clearly indicates the internal configuration in the hadron, this scaling relation can be employed to find the internal configurations of exotic hadrons. The constituent counting rule provides useful guidance for us to further explore the properties of exotic hadrons.
V Dalitz process and experimental feasibility
Figure 16: Feynman diagrams for the $\pi^{-}p\rightarrow K^{+}\pi^{-}\Sigma^{0}$ reaction.
In the $\pi^{-}p\rightarrow K^{*}\Sigma$ process, the $K^{*}$ meson produced cannot be measured directly. Instead, it must be determined by reconstructing the final particles resulting from its decay. Given that the branching ratio for $K^{*}$ decay to $K^{+}\pi^{-}$ is as high as $99.754\%$ ParticleDataGroup:2022pth , analyzing the Dalitz process for $\pi^{-}p\rightarrow K^{*}\Sigma\rightarrow K^{+}\pi^{-}\Sigma^{0}$ is crucial for providing useful information for experimental measurements. The basic tree-level Feynman diagrams for the $\pi^{-}p\rightarrow K^{+}\pi^{-}\Sigma^{0}$ reaction are illustrated in Fig. 16. Our calculations show that the total cross section for $N^{*}(2080)(\frac{1}{2}^{-})$ is slightly larger than that for the $N^{*}(2080)(\frac{3}{2}^{-})$. In this study, we focus on the Dalitz process specifically for $N^{*}$ with $J^{P}=\frac{3}{2}^{-}$. Generally, the invariant mass spectrum of the Dalitz process is defined based on the two-body process Kim:2017nxg ; Wang:2024qnk
$\frac{d\sigma_{\pi^{-}p\rightarrow K^{*}\Sigma\rightarrow K^{+}\pi^{-}\Sigma^{0}}}{dM_{K^{+}\pi^{-}}}\approx\frac{2M_{K^{*}}M_{K^{+}\pi^{-}}}{\pi}\frac{\sigma_{\pi^{-}p\rightarrow K^{*}\Sigma}\Gamma_{K^{*}\rightarrow K^{+}\pi^{-}}}{(M^{2}_{K^{+}\pi^{-}}-M^{2}_{K^{*}})^{2}+M^{2}_{K^{*}}\Gamma^{2}_{K^{*}}},$
where $\Gamma_{K^{*}}=47.3\text{ MeV}$ and $\Gamma_{K^{*}\rightarrow K^{+}\pi^{-}}=47.2\text{ MeV}$ represent the total width and the decay width of $K^{*}$ to $K^{+}\pi^{-}$, respectively. The calculation results are shown in Fig. 17, where peak values are observed at $M_{K^{+}\pi^{-}}=895\text{ MeV}$, with peaks not less than 104.57 $\mu\text{b}$/GeV. This indicates that reconstructing the $\pi^{-}p\rightarrow K^{*}\Sigma$ process through the $\pi^{-}p\rightarrow K^{+}\pi^{-}\Sigma^{0}$ process is feasible in experiments.
Figure 17: The invariant mass distribution $d\sigma_{\pi^{-}p\rightarrow K^{*}\Sigma\rightarrow K^{+}\pi^{-}\Sigma^{0}}$ /$dM_{K^{+}\pi^{-}}$ at different c.m. energies $W$ = 2.56 GeV, 2.63 GeV, 2.88 GeV, and 3.056 GeV.
The calculated total cross sections for $\pi^{-}p\rightarrow K^{*}\Sigma\rightarrow K^{+}\pi^{-}\Sigma^{0}$ and the experimental data for $\pi^{-}p\rightarrow K^{+}\pi^{-}\Sigma^{0}$ are presented in Table 6. To assess the feasibility of experimentally detecting the $K^{*}$ via $\pi^{-}p$ scattering, we calculate the ratio $\sigma(\pi^{-}p\rightarrow K^{*}\Sigma\rightarrow K^{+}\pi^{-}\Sigma^{0})/\sigma(\pi^{-}p\rightarrow K^{+}\pi^{-}\Sigma^{0})$. The minimum ratio reaches 29.67% at center-of-mass energies between 2.56 and 2.88 GeV, indicating that the total cross section and event number for the $\pi^{-}p\rightarrow K^{*}\Sigma$ process reconstructed via $\pi^{-}p\rightarrow K^{+}\pi^{-}\Sigma^{0}$ are sufficient for experimental measurements. Additionally, the ratio can reach up to 76.61% at a center-of-mass energy of 2.56 GeV, with the $s$-channel $N^{*}(2080)$ exchange contributing 73.19%, approximately 30.28 $\mu\text{b}$. These results suggest that a significant number of the $N^{*}(2080)$ can be produced in the $\pi^{-}p\rightarrow K^{*}\Sigma\rightarrow K^{+}\pi^{-}\Sigma^{0}$ process, providing strong support for future experimental efforts.
Table 6: The total cross section of $\pi^{-}p\rightarrow K^{+}\pi^{-}\Sigma^{0}$ and $\pi^{-}p\rightarrow K^{*}\Sigma\rightarrow K^{+}\pi^{-}\Sigma^{0}$. Here, the $\sigma_{1}$ represents experimental cross section datas of $\pi^{-}p\rightarrow K^{+}\pi^{-}\Sigma^{0}$ Dahl:1967pg ; Abramovich:1972rq , while $\sigma_{2}$ is the calculated cross section of $\pi^{-}p\rightarrow K^{*}\Sigma\rightarrow K^{+}\pi^{-}\Sigma^{0}$.
$W$ $(GeV)$ $\sigma_{1}$ $(\mu b)$ $\sigma_{2}$ $(\mu b)$ ($\sigma_{2}/\sigma_{1}$)
2.56 54 41.37 76.61%
2.60 46 33.93 73.76%
2.63 41 29.53 72.02%
2.86 43 12.76 29.67%
2.88 40 12.05 30.12%
As a leading hadron experimental facility, J-PARC currently provides the highest-intensity meson beams (both $K$ and $\pi$ mesons) in the several GeV energy range, making it an outstanding platform for particle and nuclear physics research Aoki:2021cqa . Theoretical studies and experimental evidence suggest that certain light-flavored baryons may contain a significant strange quark pair ($s\bar{s}$) component, potentially forming a hidden-strange pentaquark state, such as the $N^{*}(2080)$. However, the existence of a hidden-strange pentaquark state has not yet been confirmed experimentally, underscoring the need for collaborative efforts between experimental and theoretical research. If J-PARC can measure the $\pi^{-}p\rightarrow K^{*}\Sigma$ process, the findings could aid in the discovery of the hidden-strange pentaquark state $N^{*}(2080)$ and provide valuable insights into quark interactions within hadrons, particularly the internal dynamics of diquarks.
Additionally, HIAF offers high-intensity proton and ion beam currents across a wide energy range. Upon completion, HIAF will deliver the highest peak-intensity medium and low-energy heavy-ion beam currents compared to similar facilities worldwide HIAF . If a meson beam experimental device is developed at HIAF in the future, it could enable high-precision measurements of the $\pi^{-}p\rightarrow K^{*}\Sigma$ process, further supporting the search for the $N^{*}(2080)$ wl:2024ext .
VI Summary and prospect
In recent years, the $N^{*}(2080)$ has been considered the hidden-strange partner of the hidden-charm molecular pentaquark $P_{c}(4457)$, as its mass lies just below the $K^{*}\Sigma$ threshold He:2017aps ; Ben:2023uev ; Xie:2010yk ; Lin:2018kcc . In this study, we investigate the production of the nucleon excited state $N^{*}(2080)$ through the $\pi^{-}p$ scattering process using the effective Lagrangian approach. Specifically, we consider $K$ and $K^{*}$ meson exchanges in the $t$-channel, $\Sigma$ exchange in the $u$-channel, and the contribution of the $N^{*}(2080)$ in the $s$-channel. By fitting the total and differential cross sections of the $\pi^{-}p\rightarrow K^{*}\Sigma$ process, we demonstrate that the $N^{*}(2080)$ exchange provides the dominant contribution to the total cross section near the threshold energy region, offering potential insights into detecting the $N^{*}(2080)$ through the $\pi^{-}p$ scattering process. However, the lack of experimental data near the threshold imposes limitations on determining the properties of the $N^{*}(2080)$ via this process.
Given the significant branching ratio of $K^{*}$ decay into $K^{+}\pi^{-}$, we calculate the Dalitz process for $\pi^{-}p\rightarrow K^{*}\Sigma\rightarrow K^{+}\pi^{-}\Sigma^{0}$ and compare it with experimental data for $\pi^{-}p\rightarrow K^{+}\pi^{-}\Sigma^{0}$. The results suggest that it is feasible to observe the $N^{*}(2080)$ through the $\pi^{-}p\rightarrow K^{*}\Sigma$ process in experiments. Notably, our calculations consider two scenarios for the $J^{P}$ quantum number of the $N^{*}(2080)$, either $\frac{1}{2}^{-}$ or $\frac{3}{2}^{-}$. A significant source of uncertainty in our model arises from the absence of experimental measurements of the partial widths of the $N^{*}(2080)$ decays to $\pi N$, leading us to treat the coupling constants $g_{\pi NN^{*}}^{3/2^{-}}$ and $g_{\pi NN^{*}}^{1/2^{-}}$ as free parameters determined by fitting. In conclusion, further experimental measurements and theoretical investigations into the internal properties of $N^{*}$ states are highly desirable.
Additionally, the application of the constituent counting rule in hard exclusive reactions is discussed, including a demonstration of the rule by deriving expressions for the reaction cross-section and matrix element, as well as explaining the contribution of typical hard gluon exchange processes to exclusive reactions. By calculating the mass dimension, we derive the constituent counting expression for the reaction cross-section. In the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction, the predictions of this rule align with experimental data, further validating its accuracy.
Currently, high-precision measurements of the meson-nucleus scattering process can be conducted by experiments at J-PARC Aoki:2021cqa , AMBER Adams:2018pwt , and future facilities such as HIKE HIKE:2023ext and HIAF HIAF . We recommend that these experiments perform precise measurements of the $\pi^{-}p\rightarrow K^{*}\Sigma$ process, particularly the differential cross section at the forward angle ($\theta=0^{\circ}$) and the $t$-distribution cross section at large momentum transfer ($\theta=90^{\circ}$). This is crucial for elucidating the mechanism of the $\pi^{-}p\rightarrow K^{*}\Sigma$ process and determining the contribution of the $N^{*}$ state. To date, pentaquark-like states containing strange quarks have not been experimentally confirmed, necessitating further clarification through collaborative experimental and theoretical efforts. Analyzing the existence and contribution of pentaquark-like states with strange quarks via the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction is an effective approach wl:2024ext . Moving forward, we plan to systematically investigate the contribution of excited baryon states in various $\pi^{-}p$ or $K^{-}p$ scattering processes, providing essential theoretical support for future experiments.
This study is part of our ongoing efforts. The experimental beam currents and related data necessary for specific hadron spectroscopy studies can only be determined through Monte Carlo simulations based on the parameters (such as luminosity, etc.) of experimental facilities like HIAF. We will also collaborate closely with experimentalists to conduct further in-depth research in this area.
VII Acknowledgments
This work is supported by the National Natural Science Foundation of China under Grants No. 12065014, No. 12047501 and No. 12247101, the Natural Science Foundation of Gansu province under Grant No. 22JR5RA266, and the West Light Foundation of The Chinese Academy of Sciences under Grant No. 21JR7RA201. X.L. is also supported by National Natural Science Foundation of China under Grant No. 12335001, National Key Research and Development Program of China under Contract No. 2020YFA0406400, the 111 Project under Grant No. B20063, the Fundamental Research Funds for the Central Universities, and the project for top-notch innovative talents of Gansu province.

### v2

@@
-Thorough Analysis of $\pi^{-}p\rightarrow K^{*}\Sigma$ Reaction: Probing Nucleon Excited State and Constituent Scaling Laws
+Production potential of hidden-strange molecular pentaquarks through the $\pi^{-}p\rightarrow K^{*}\Sigma$ process
 Abstract
-In this study, applying the effective Lagrangian approach, we investigate the reaction $\pi^{-}p\rightarrow K^{*}\Sigma$ to explore the production of the nucleon excited state $N^{*}(2080)$ and the constituent field scaling laws, also known as the constituent counting rules.
-Specifically, we analyze two scenarios, where the $J^{P}$ quantum numbers of $N^{*}(2080)$ are $3/2^{-}$ and $1/2^{-}$.
-Our results show that the contribution from $s$-channel $N^{*}$ exchange dominates the total cross section, surpassing those from $t$-channel $K^{(*)}$ exchange and $u$-channel $\Sigma$ exchange.
+In this study, applying the effective Lagrangian approach, we investigate the reaction $\pi^{-}p\rightarrow K^{*}\Sigma$ to explore the production of the hidden-strange molecular pentaquarks $P_{s\bar{s}}$.
+Specifically, we analyze two scenarios, where the $J^{P}$ quantum numbers of $P_{s\bar{s}}$ are $3/2^{-}$ and $1/2^{-}$.
+Our results show that the contribution from $s$-channel $P_{s\bar{s}}$ exchange dominates the total cross section, surpassing those from $t$-channel $K^{(*)}$ exchange and $u$-channel $\Sigma$ exchange.
 Additionally, we present predictions for the differential cross section of the $\pi^{-}p\rightarrow K^{*}\Sigma$ process.
-Notably, in the large momentum transfer region ($\theta=90^{\circ}$), the differential cross section in the $t$-distribution aligns with the energy-dependent scaling formula predicted by the constituent counting rule, specifically, $d\sigma/dt\sim 1/s^{8}$ for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction.
 Finally, we extend the 2-to-2 scattering process to the 2-to-3 Dalitz process and provide theoretical predictions for this scenario.
 Our findings suggest that the cross section for the $\pi^{-}p\rightarrow K^{*}\Sigma\rightarrow K^{+}\pi^{-}\Sigma^{0}$ process can reach several tens of $\mu b$ near the threshold, making it highly favorable for experimental measurement.
-However, the current scarcity of experimental data for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction at threshold energy limits our ability to precisely determine the properties of the nucleon excited state $N^{*}$ and related physical quantities.
+However, the current scarcity of experimental data for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction at threshold energy limits our ability to determine the properties of the hidden-strange molecular pentaquarks $P_{s\bar{s}}$ and related physical quantities.
 Therefore, additional experimental measurements of the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction at threshold are strongly encouraged, and correlation studies could be pursued at facilities such as J-PARC, AMBER, and upcoming HIKE and HIAF meson beam experiments.
 I INTRODUCTION
-Over the past several decades, meson-nucleon scattering experiments have played a crucial role in the development of hadron spectroscopy wl:2024ext .
-These experiments are particularly significant, as more than $80\%$ of the light hadrons cataloged by the Particle Data Group (PDG) ParticleDataGroup:2022pth have been discovered or characterized through such studies.
-Hadron spectroscopy provides an effective means to enhance our understanding of the nonperturbative aspects of strong interactions.
-In recent years, theorists have also focused on scattering processes involving $K$ and $\pi$ mesons, aiming to explore the production of new hadronic states Liu:2019zoy ; Chen:2016qju ; Chen:2022asf ; Zou:2016bxw ; Lin:2022eau ; Wang:2023lia ; Wang:2022sib ; Xiang:2020phx ; Wang:2019zaw ; Wang:2019uwk ; Wang:2019dsi ; Wang:2018mjz ; Wang:2017qcw ; Cheng:2016ddp ; Wang:2015xwa ; Liu:2021ojf ; Gao:2012zh ; Wu:2009tu ; Xie:2014kja ; Xie:2007qt .
-Both experimental and theoretical studies suggest that meson beam experiments offer unique advantages in advancing research on hadron spectroscopy and related physical phenomena.
-Light-flavor baryons have garnered significant attention from the hadron physics community due to the well-known “missing” problem associated with their excited states ZOU:2005 .
+Hadron spectroscopy offers an effective approach to deepen our understanding of the nonperturbative aspects of strong interactions.
+Over the past several decades, meson-nucleon scattering experiments have played a crucial role in the advancement of hadron spectroscopy wl:2024ext .
+This is evident from the fact that over $80\%$ of the light hadrons cataloged by the Particle Data Group (PDG) ParticleDataGroup:2022pth have been discovered through such scattering processes.
+While collider experiments are the primary means of exploring the nature of matter, scattering experiments deserve greater attention.
+In recent years, theorists have increasingly focused on scattering processes involving $K$ and $\pi$ meson beams, aiming to investigate the production of new hadronic states Liu:2019zoy ; Chen:2016qju ; Chen:2022asf ; Zou:2016bxw ; Lin:2022eau ; Wang:2023lia ; Wang:2022sib ; Xiang:2020phx ; Wang:2019zaw ; Wang:2019uwk ; Wang:2019dsi ; Wang:2018mjz ; Wang:2017qcw ; Cheng:2016ddp ; Wang:2015xwa ; Liu:2021ojf ; Gao:2012zh ; Wu:2009tu ; Xie:2014kja ; Xie:2007qt .
+With the observation of several $P_{c}$ states by the LHCb Collaboration LHCb:2015yax ; LHCb:2019kea , hidden-charm molecular-type pentaquarks, predicted in Refs.
+Chen:2015loa ; Chen:2015moa ; Wang:2019krd ; Shimizu:2016rrd ; Roca:2015dva ; Wang:2019dsi , have been identified.
+This identification stems from the discovery of a characteristic mass spectrum in the $J/\psi p$ invariant mass distribution from the decay $\Lambda_{b}\to J/\psi pK$ LHCb:2015yax ; LHCb:2019kea .
+This represents an important progress in the study of hadron spectroscopy.
+Following this, it is natural to further investigate the hidden-strange partners of these hidden-charm molecular-type pentaquarks while developing the “Particle Zoo 2.0” version.
 Among the numerous observed light-flavor baryon states, the $N^{*}(2080)$ stands out as a particularly intriguing research target, closely linked to the investigation of hidden strange pentaquark states.
@@
 The potential existence of such hadronic molecules has also been suggested in studies utilizing the quark delocalization color screening model Gao:2017hya .
-As discussed in Ref.
-wl:2024ext , hidden strange pentaquark states could be identified and studied by measuring processes such as $\pi^{-}p\rightarrow K^{(*)}\Sigma^{(*)}$ and $\pi^{-}p\rightarrow K^{(*)}\Lambda^{(*)}$.
-These measurements could provide crucial experimental data to address the “missing” problem of baryon excited states.
 In Ref.
@@
 The results showed that the measured decay properties of $N^{*}(2080)$ with $J^{P}=3/2^{-}$ were well reproduced, supporting the molecular interpretation of this state.
-Given that the $J^{P}$ quantum numbers of the $N^{*}(2080)$ [^1] According to the latest PDG ParticleDataGroup:2022pth , the previously identified two-star nucleon resonance $N^{*}(2080)$ has been split into a three-star $N(1875)$ and a two-star $N(2120)$, both with spin-parity $3/2^{-}$.
-A key characteristic of hadronic molecules is that they are typically very shallowly bound Guo:2017jvc , meaning that their masses are generally just below the corresponding threshold.
-In this study, we consider the $K^{*}\Sigma$ molecular interpretation, with a threshold of 2086 MeV, and accordingly, we adopt the mass of the $N(2120)$ as 2080 MeV, referring to it by its previous designation, the $N^{*}(2080)$, throughout the remainder of this paper.
-could be $1/2^{-}$ or $3/2^{-}$, this state can couple to $K^{*}\Sigma$ through $S$ wave interaction.
-The coupling constant between $N^{*}$ and $K^{*}\Sigma$ can be determined using the Weinberg compositeness criterion Lin:2018kcc ; Weinberg:1965zz ; Baru:2003qq ; Lin:2017mtz .
-Therefore, searching for and measuring the $N^{*}(2080)$ state via the $\pi^{-}p\rightarrow K^{*}\Sigma$ process is a feasible approach.
-In the 1970s and 1980s, several experiments measured scattering cross section data for the $\pi^{-}p\rightarrow K^{*}\Sigma$ process Crennell:1972km ; CERN-CollegedeFrance-Madrid-Stockholm:1980mch ; CERN-CollegedeFrance-Madrid-Stockholm:1980ysu ; Dahl:1967pg ; Miller:1965 ; Abramovich:1972rq , providing crucial support for our research.
-In this work, we will perform calculations and analysis of the $\pi^{-}p\rightarrow K^{*}\Sigma$ scattering process using the effective Lagrangian approach, with the aim of investigating the production of the nucleon excited state $N^{*}(2080)$ and exploring the constituent counting rule.
+We observe a change in the resonance designation relation to $N^{*}(2080)$.
+According to the latest PDG ParticleDataGroup:2022pth , the previously identified two-star nucleon resonance $N^{*}(2080)$ has been split into a three-star $N(1875)$ and a two-star $N(2120)$, both with spin-parity $3/2^{-}$.
+Therefore, in our analysis, we no longer use $N^{*}(2080)$ to represent the hidden-strange molecular-type pentaquarks as considered in Refs.
+He:2017aps ; Lin:2018kcc ; Khemchandani:2011et ; Doring:2008sv .
+Instead, we focus on the production of hidden-strange molecular-type pentaquarks with the $K^{*}\Sigma$ component.
+This leads to two possible $J^{P}$ quantum numbers for S-wave $K^{*}\Sigma$ molecular systems: $J^{P}=1/2^{-}$ and $3/2^{-}$.
+In this study, we designate two S-wave $K^{*}\Sigma$ molecular-type pentaquarks with $J^{P}=1/2^{-}$ and $3/2^{-}$ as $P_{s\bar{s}}[1/2^{-}]$ and $P_{s\bar{s}}[3/2^{-}]$, respectively.
+In this work, we investigate the production of two possible hidden-strange molecular-type pentaquarks through the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction.
+We notice several previous experimental results, where the measured scattering cross section data for the $\pi^{-}p\rightarrow K^{*}\Sigma$ process were given Crennell:1972km ; CERN-CollegedeFrance-Madrid-Stockholm:1980mch ; CERN-CollegedeFrance-Madrid-Stockholm:1980ysu ; Dahl:1967pg ; Miller:1965 ; Abramovich:1972rq .
+It provides a good chance to combining with these experimental data to carry out the present research.
+For quantitatively calculating the $\pi^{-}p\rightarrow K^{*}\Sigma$ scattering process, we adopt the effective Lagrangian approach, with the aim of investigating the production of hidden-strange molecular pentaquarks $P_{s\bar{s}}$.
 Additionally, we will analyze the $\pi^{-}p\rightarrow K^{*}\Sigma\rightarrow K^{+}\pi^{-}\Sigma^{0}$ Dalitz process and assess its experimental feasibility, thereby offering valuable theoretical insights for future experimental studies.
-Actually, the $\pi^{-}p\rightarrow K^{*}\Sigma$ process serves as an ideal case for examining the constituent field scaling law, commonly known as the constituent counting rule.
-Initially proposed by Brodsky in 1973 and further developed in the decade following the advent of perturbative QCD (pQCD), this rule embodies the principles of conformality and scale invariance in QCD at high energies Lepage:1980fj ; Mueller:1981sg ; Lepage:1979zb ; Brodsky:1974vy ; Brodsky:1973kr ; Matveev:1973ra ; Guo:2016fqg .
-The constituent counting rule has been nonperturbatively derived in contexts like AdS/QCD, showcasing its broad applicability to various field theories Polchinski:2001tt .
-In essence, the rule predicts that in exclusive processes such as $a+b\rightarrow c+d$ involving large momentum transfers, quarks must share significant momenta to coalesce into hadrons by exchanging hard gluons.
-Consequently, the cross-section for such an exclusive reaction scales according to $d\sigma/dt\sim s^{2-n}f(\theta_{\mathrm{c.m.}})$, where $n=n_{a}+n_{b}+n_{c}+n_{d}$ represents the total number of constituents in particles $a$, $b$, $c$, and $d$.
-Here, $s$ and $t$ are the Mandelstam variables, and $\theta_{\mathrm{c.m.}}$ is the scattering angle in the center-of-mass frame.
-The constituent counting rule has been validated by two-body hadronic exclusive reactions conducted at the Brookhaven National Laboratory’s Alternating Gradient Synchrotron (BNL-AGS) Baller:1988tj .
-More broadly, this rule effectively describes the differential cross-section behavior in various hadron-hadron scattering processes within the large momentum transfer regime JeffersonLabHallA:2002vjd ; JeffersonLabHallA:2004gxp ; White:1994tj ; Amaryan:2021cnj .
-Applying the constituent counting rule in our analysis of the $\pi^{-}p\rightarrow K^{*}\Sigma$ process not only deepens our understanding of the mechanisms underlying hadron-hadron scattering and the role of constituent quarks but also allows us to assess the accuracy of our theoretical cross-section predictions.
 This paper is organized as follows.
 After the introduction, the Lagrangians and amplitudes used in this work are presented in Section II.
-The numerical results for the total and differential cross sections are provided in Section III.
-In Section IV, we introduce the constituent counting rule and assess its validity.
-The Dalitz process and its experimental feasibility are discussed in Section V, followed by a brief summary in Section VI.
-II THE PRODUCTION OF THE $N^{*}(2080)$ VIA THE $\pi^{-}p\rightarrow K^{*}\Sigma$ REACTION
+The numerical results for the cross sections and the constituent counting rule are provided in Section III.
+The Dalitz process and its experimental feasibility are discussed in Section IV, followed by a brief summary in Section V.
+II THE PRODUCTION OF THE $P_{s\bar{s}}$ VIA THE $\pi^{-}p\rightarrow K^{*}\Sigma$ REACTION
 The schematic tree-level Feynman diagrams for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction are depicted in Fig.
 1.
-These include $s$-channel $N^{*}(2080)$ ($\equiv N^{*}$) exchange, $u$-channel $\Sigma$ exchange, and $t$-channel $K$ and $K^{*}$ exchanges.
+These include $s$-channel $P_{s\bar{s}}$ ($\equiv P^{*}$) exchange, $u$-channel $\Sigma$ exchange, and $t$-channel $K$ and $K^{*}$ exchanges.
 Figure 1: Feynman diagrams for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction.
 To evaluate the pion-induced production of $K^{*}\Sigma$, the relevant Lagrangians are required, as used in previous studies Xiang:2020phx ; Kim:2015ita ; He:2017aps ; Xie:2014kja ; Ben:2023uev ; Wang:2019dsi ; Cheng:2016ddp .
-In general, similar to the case of the $P_{c}$ pentaquark state, the $J^{P}$ quantum numbers of the $N^{*}(2080)$ state can be either $\frac{3}{2}^{-}$ or $\frac{1}{2}^{-}$.
-Therefore, for the $s$-channel, both quantum number assignments for the $N^{*}(2080)$ should be considered in our calculations.
+In general, similar to the case of the $P_{c}$ pentaquark state, the $J^{P}$ quantum numbers of the $P_{s\bar{s}}$ state can be either $\frac{3}{2}^{-}$ or $\frac{1}{2}^{-}$.
+Therefore, for the $s$-channel, both quantum number assignments for the $P_{s\bar{s}}$ should be considered in our calculations.
 These Lagrangians include
-$\mathcal{L}^{3/2^{-}}_{K^{*}\Sigma N^{*}}$ $=$ $g_{K^{*}\Sigma N^{*}}^{3/2^{-}}\bar{N}^{*}_{\mu}{\Sigma}K^{*\mu}+\mathrm{H.c.},$ (1)
-$\mathcal{L}^{3/2^{-}}_{\pi NN^{*}}$ $=$ $\frac{g_{\pi NN^{*}}^{3/2^{-}}}{m_{\pi}}\bar{N}\gamma_{5}\bm{\tau}\cdot\partial^{\mu}\bm{\pi}N^{*}_{\mu}+\mathrm{H.c.},$ (2)
-$\mathcal{L}^{1/2^{-}}_{K^{*}\Sigma N^{*}}$ $=$ $g_{K^{*}\Sigma N^{*}}^{1/2^{-}}\bar{N}^{*}\gamma_{5}\gamma^{\mu}K^{*}_{\mu}\Sigma+\mathrm{H.c.},$ (3)
-$\mathcal{L}^{1/2^{-}}_{\pi NN^{*}}$ $=$ $g_{\pi NN^{*}}^{1/2^{-}}\bar{N}\bm{\tau}\cdot\bm{\pi}N^{*}+\mathrm{H.c.}.$ (4)
+$\mathcal{L}^{3/2^{-}}_{K^{*}\Sigma P^{*}}$ $=$ $g_{K^{*}\Sigma P^{*}}^{3/2^{-}}\bar{P^{*}}_{\mu}{\Sigma}K^{*\mu}+\mathrm{H.c.},$ (1)
+$\mathcal{L}^{3/2^{-}}_{\pi NP^{*}}$ $=$ $\frac{g_{\pi NP^{*}}^{3/2^{-}}}{m_{\pi}}\bar{N}\gamma_{5}\bm{\tau}\cdot\partial^{\mu}\bm{\pi}P^{*}_{\mu}+\mathrm{H.c.},$ (2)
+$\mathcal{L}^{1/2^{-}}_{K^{*}\Sigma P^{*}}$ $=$ $g_{K^{*}\Sigma P^{*}}^{1/2^{-}}\bar{P^{*}}\gamma_{5}\gamma^{\mu}K^{*}_{\mu}\Sigma+\mathrm{H.c.},$ (3)
+$\mathcal{L}^{1/2^{-}}_{\pi NP^{*}}$ $=$ $g_{\pi NP^{*}}^{1/2^{-}}\bar{N}\bm{\tau}\cdot\bm{\pi}P^{*}+\mathrm{H.c.}.$ (4)
 The Lagrangians corresponding to the $u$-channel are written as
@@
 The symbols $\bm{\pi}$, $K$, and $K^{*}$ represent the $\pi$, $K$, and $K^{*}$ meson fields, respectively.
-Similarly, $N$, $\Sigma$, and $N^{*}$ correspond to the nucleon, $\Sigma$ baryon, and $N^{*}(2080)$ fields, respectively.
-Assuming the $N^{*}(2080)(\frac{3}{2}^{-})$ is a pure $S$-wave molecular state of $K^{*}$ and $\Sigma$, the coupling constant $g_{K^{*}\Sigma N^{*}}^{3/2^{-}}$ can be estimated model-independently using the Weinberg compositeness criterion, yielding Weinberg:1965zz ; Baru:2003qq ; Lin:2017mtz ; Lin:2018kcc
-$g_{K^{*}\Sigma N^{*}}^{2}$ $=$ $\frac{4\pi}{4M_{N^{*}}M_{\Sigma}}\frac{(M_{K^{*}}+M_{\Sigma})^{5/2}}{(M_{K^{*}}M_{\Sigma})^{1/2}}\sqrt{32\epsilon},$ (11)
-$\epsilon$$\equiv$ $M_{K^{*}}+M_{\Sigma}-M_{N^{*}},$ (12)
-where $M_{N^{*}}$, $M_{K^{*}}$, and $M_{\Sigma}$ represent the masses of the $N^{*}(2080)(\frac{3}{2}^{-})$, $K^{*}$, and $\Sigma$, respectively, while $\epsilon$ denotes the $K^{*}\Sigma$ binding energy.
+Similarly, $N$, $\Sigma$, and $P^{*}$ correspond to the nucleon, $\Sigma$ baryon, and $P_{s\bar{s}}$ fields, respectively.
+Assuming the $P_{s\bar{s}}[3/2^{-}]$ is a pure $S$-wave molecular state of $K^{*}$ and $\Sigma$, the coupling constant $g_{K^{*}\Sigma P^{*}}^{3/2^{-}}$ can be estimated model-independently using the Weinberg compositeness criterion, yielding Weinberg:1965zz ; Baru:2003qq ; Lin:2017mtz ; Lin:2018kcc
+$g_{K^{*}\Sigma P^{*}}^{2}$ $=$ $\frac{4\pi}{4M_{P^{*}}M_{\Sigma}}\frac{(M_{K^{*}}+M_{\Sigma})^{5/2}}{(M_{K^{*}}M_{\Sigma})^{1/2}}\sqrt{32\epsilon},$ (11)
+$\epsilon$$\equiv$ $M_{K^{*}}+M_{\Sigma}-M_{P^{*}},$ (12)
+where $M_{P^{*}}$, $M_{K^{*}}$, and $M_{\Sigma}$ represent the masses of the $P_{s\bar{s}}[3/2^{-}]$, $K^{*}$, and $\Sigma$, respectively, while $\epsilon$ denotes the $K^{*}\Sigma$ binding energy.
 Assuming the physical state is a pure $S$-wave hadronic molecule, the relative uncertainty of the above approximation for the coupling constant is given by $\sqrt{2\mu\epsilon}\cdot r$, where $\mu=\frac{m_{1}m_{2}}{m_{1}+m_{2}}$ is the reduced mass of the bound particles, and $r$ is the range of forces, which can be estimated as the inverse of the mass of the exchanged particle.
 For the $K^{*}\Sigma$ system, $r$ can be estimated as $1/m_{K}$ Lin:2018kcc .
-Following Ref. Lin:2018kcc , we adopt the mass of the $N^{*}(2080)(\frac{3}{2}^{-})$ as $M_{N^{*}}=2080\text{ MeV}$. Substituting this into Eq. (11), one obtains:
-$g_{K^{*}\Sigma N^{*}}^{3/2^{-}}=1.72.$ (13)
-Since the $N^{*}(2080)(\frac{1}{2}^{-})$ can also be considered a $S$-wave molecular state of $K^{*}$ and $\Sigma$, we approximate the value of $g_{K^{*}\Sigma N^{*}}^{1/2^{-}}$ as 1.72.
-In the absence of experimental measurements for the partial width of the $N^{*}(2080)$ decays to $\pi N$, we treat the coupling constants $g_{\pi NN^{*}}^{3/2^{-}}$ and $g_{\pi NN^{*}}^{1/2^{-}}$ as free parameters to be determined through fitting.
+Following Ref. Lin:2018kcc , we adopt the mass of the $P_{s\bar{s}}[3/2^{-}]$ as $M_{P^{*}}=2080\text{ MeV}$. Substituting this into Eq. (11), one obtains:
+$g_{K^{*}\Sigma P^{*}}^{3/2^{-}}=1.72.$ (13)
+Since the $P_{s\bar{s}}[1/2^{-}]$ can also be considered a $S$-wave molecular state of $K^{*}$ and $\Sigma$, we approximate the value of $g_{K^{*}\Sigma P^{*}}^{1/2^{-}}$ as 1.72.
+In the absence of experimental measurements for the partial width of the $P_{s\bar{s}}$ decays to $\pi N$, we treat the coupling constants $g_{\pi NP^{*}}^{3/2^{-}}$ and $g_{\pi NP^{*}}^{1/2^{-}}$ as free parameters to be determined through fitting.
 The coupling constant values for the $\Sigma\Sigma\pi$, $K^{*}N\Sigma$, and $KN\Sigma$ interactions have been provided in several theoretical works Xiang:2020phx ; He:2017aps ; Cheng:2016ddp .
@@
 For the $t$-channel and $u$-channel, the form factors are adopted in our calculation as follows,
-$F_{t/u}(q)=\frac{\Lambda{{}^{2}_{t/u}}-m^{2}}{\Lambda{{}^{2}_{t/u}}-q^{2}}.$ (16)
+$F_{t/u}(q)=\left(\frac{\Lambda{{}^{2}_{t/u}}-m^{2}}{\Lambda{{}^{2}_{t/u}}-q^{2}}\right)^{2}.$ (16)
 To account for the finite size of the hadron, we incorporate a form factor in our calculation for the $s$-channel with an intermediate baryon, as follows:
@@
 The reduced amplitudes $\mathcal{A}_{s,\mu}$, $\mathcal{A}_{u,\mu}$, and $\mathcal{A}_{t,\mu}$ for the $s$-, $u$-, and $t$-channel contributions read as
-$\mathcal{A}^{N^{*({3/2}^{-})}}_{s,\mu}$ $=$ $-i\sqrt{2}g_{K^{*}\Sigma{N^{*}}}^{{3/2}^{-}}\frac{g_{\pi N{N^{*}}}^{{3/2}^{-}}}{m_{\pi}}F_{s}(q)\gamma_{5}k_{1}^{\nu}$ (19)
-×$\times\frac{(q\mkern-8.0mu/_{s}+m_{N^{*}})}{s-m^{2}_{N^{*}}+im_{N^{*}}\Gamma_{N^{*}}}\Delta_{\mu\nu},$
-$\mathcal{A}^{N^{*({1/2}^{-})}}_{s,\mu}$ $=$ $-\sqrt{2}g_{K^{*}\Sigma{N^{*}}}^{{1/2}^{-}}g_{\pi N{N^{*}}}^{{1/2}^{-}}F_{s}(q)\gamma_{5}\gamma_{\mu}$ (20)
-×$\times\frac{(q\mkern-8.0mu/_{s}+m_{N^{*}})}{s-m^{2}_{N^{*}}+im_{N^{*}}\Gamma_{N^{*}}},$
-$\mathcal{A}^{\Sigma}_{u,\mu}$ $=$ $-\sqrt{2}\frac{g_{\Sigma\Sigma\pi}}{m_{\pi}}g_{K^{*}N\Sigma}F_{u}(q)\gamma^{5}\gamma^{\nu}k_{1\nu}\frac{(q\mkern-8.0mu/_{\Sigma}+m_{\Sigma})}{{u-m_{\Sigma}^{2}}}$ (21)
+$\mathcal{A}^{P^{*({3/2}^{-})}}_{s,\mu}$ $=$ $\sqrt{2}g_{K^{*}\Sigma P^{*}}^{{3/2}^{-}}\frac{g_{\pi NP^{*}}^{{3/2}^{-}}}{m_{\pi}}F_{s}(q)\gamma_{5}k_{1}^{\nu}$ (19)
+×$\times\frac{(q\mkern-8.0mu/_{s}+m_{P^{*}})}{s-m^{2}_{P^{*}}+im_{P^{*}}\Gamma_{P^{*}}}\Delta_{\mu\nu},$
+$\mathcal{A}^{P^{*({1/2}^{-})}}_{s,\mu}$ $=$ $-i\sqrt{2}g_{K^{*}\Sigma P^{*}}^{{1/2}^{-}}g_{\pi NP^{*}}^{{1/2}^{-}}F_{s}(q)\gamma_{5}\gamma_{\mu}$ (20)
+×$\times\frac{(q\mkern-8.0mu/_{s}+m_{P^{*}})}{s-m^{2}_{P^{*}}+im_{P^{*}}\Gamma_{P^{*}}},$
+$\mathcal{A}^{\Sigma}_{u,\mu}$ $=$ $-i\sqrt{2}\frac{g_{\Sigma\Sigma\pi}}{m_{\pi}}g_{K^{*}N\Sigma}F_{u}(q)\gamma^{5}\gamma^{\nu}k_{1\nu}\frac{(q\mkern-8.0mu/_{\Sigma}+m_{\Sigma})}{{u-m_{\Sigma}^{2}}}$ (21)
 ×$\times(\gamma_{\mu}-\frac{k_{K^{*}{\Sigma}N}}{2m_{N}}\gamma_{\mu}q\mkern-8.0mu/_{K^{*}}),$
-$\mathcal{A}^{K}_{t,\mu}$ $=$ $-\sqrt{2}g_{K^{*}K\pi}g_{K\Sigma N}F_{t}(q)\frac{1}{{t-m_{K}^{2}}}\gamma_{5}$ (22)
+$\mathcal{A}^{K}_{t,\mu}$ $=$ $i\sqrt{2}g_{K^{*}K\pi}g_{K\Sigma N}F_{t}(q)\frac{1}{{t-m_{K}^{2}}}\gamma_{5}$ (22)
 ×$\times[k_{1\mu}+(k_{1}-k_{2})_{\mu}],$
-$\mathcal{A}^{K^{*}}_{t,\mu}$ $=$ $-i\sqrt{2}g_{\pi K^{*}K^{*}}g_{K^{*}N\Sigma}F_{t}(q)(\gamma_{\xi}-\frac{k_{K^{*}{\Sigma}N}}{2m_{N}}\gamma_{\xi}q\mkern-8.0mu/_{K^{*}})$ (23)
+$\mathcal{A}^{K^{*}}_{t,\mu}$ $=$ $i\sqrt{2}g_{\pi K^{*}K^{*}}g_{K^{*}N\Sigma}F_{t}(q)(\gamma_{\xi}-\frac{k_{K^{*}{\Sigma}N}}{2m_{N}}\gamma_{\xi}q\mkern-8.0mu/_{K^{*}})$ (23)
 ×$\times\epsilon_{\mu\nu\alpha\beta}\frac{\mathcal{P}^{\nu\xi}}{{t-m_{K^{*}}^{2}}}k_{2}^{\alpha}(k_{1}-k_{2})^{\beta}$
@@
 $\vec{k}_{1}^{c.m.}$ and $\vec{k}_{2}^{c.m.}$ denote the three-momenta of the initial $\pi$ beam and the final $K^{*}$ meson, respectively.
-III.1 Production of the $N^{*}(2080)$ with $J^{P}$ = $\frac{3}{2}^{-}$
+III.1 Production of the $P_{s\bar{s}}$ with $J^{P}$ = $\frac{3}{2}^{-}$
 Using the MINUIT code from CERNLIB, we will fit the experimental data CERN-CollegedeFrance-Madrid-Stockholm:1980ysu ; CERN-CollegedeFrance-Madrid-Stockholm:1980mch ; Crennell:1972km ; Abramovich:1972rq ; Dahl:1967pg ; Miller:1965 for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction.
@@
 The fitting parameters are listed in Table 1.
-The fit excluding the contribution of the $N^{*}(2080)(\frac{3}{2}^{-})$ yields $\chi^{2}/d.o.f.$ = 6.19, while including the $N^{*}(2080)(\frac{3}{2}^{-})$ improves the fit to $\chi^{2}/d.o.f.$ = 1.52.
-This indicates a significant contribution from the $s$-channel $N^{*}(2080)(\frac{3}{2}^{-})$.
+The fit excluding the contribution of the $P_{s\bar{s}}[3/2^{-}]$ yields $\chi^{2}/d.o.f.$ = 6.19, while including the $P_{s\bar{s}}[3/2^{-}]$ improves the fit to $\chi^{2}/d.o.f.$ = 1.52.
+This indicates a significant contribution from the $s$-channel $P_{s\bar{s}}[3/2^{-}]$.
 Table 1: Fitted values of the free parameters with all experimental data in Ref.
-CERN-CollegedeFrance-Madrid-Stockholm:1980ysu ; CERN-CollegedeFrance-Madrid-Stockholm:1980mch ; Crennell:1972km ; Abramovich:1972rq ; Dahl:1967pg ; Miller:1965 for the case of the $N^{*}(2080)$ with $J^{P}=\frac{3}{2}^{-}$.
-with $N^{*}(2080)(\frac{3}{2}^{-})$ without $N^{*}(2080)(\frac{3}{2}^{-})$
+CERN-CollegedeFrance-Madrid-Stockholm:1980ysu ; CERN-CollegedeFrance-Madrid-Stockholm:1980mch ; Crennell:1972km ; Abramovich:1972rq ; Dahl:1967pg ; Miller:1965 for the case of the $P_{s\bar{s}}$ with $J^{P}=\frac{3}{2}^{-}$.
+with $P_{s\bar{s}}[3/2^{-}]$ without $P_{s\bar{s}}[3/2^{-}]$
 $\Lambda_{K}$ 1.53$\pm$0.04 1.90$\pm$0.01
@@
 $\Lambda_{s}$ 1.18$\pm$0.06
-$g_{\pi N{N^{*}}}^{{3/2}^{-}}$ 0.37$\pm$0.04
+$g_{\pi N{P^{*}}}^{{3/2}^{-}}$ 0.37$\pm$0.04
 $\chi^{2}/d.o.f.$ 1.52 6.19
-The total cross section for the reaction $\pi^{-}p\rightarrow K^{*}\Sigma$ with the $N^{*}(2080)$ having $J^{P}$ = $\frac{3}{2}^{-}$ is illustrated in Fig.
+The total cross section for the reaction $\pi^{-}p\rightarrow K^{*}\Sigma$ with the $P_{s\bar{s}}$ having $J^{P}$ = $\frac{3}{2}^{-}$ is illustrated in Fig.
 2.
@@
 The solid (red), dashed-dotted (blue), dashed (purple), dash–double dotted (green) lines are for the full model, the $s$-channel, the $t$-channel and the $u$-channel, respectively.
-Here, the spin-parity quantum number of the $N^{*}(2080)$ is $\frac{3}{2}^{-}$.
+Here, the spin-parity quantum number of the $P_{s\bar{s}}$ is $\frac{3}{2}^{-}$.
 Figure 3: The differential cross-section $d\sigma/dcos\theta$ for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction as a function of $\cos\theta$.
@@
 In Fig.
-3, we present the differential cross section for the reaction $\pi^{-}p\rightarrow K^{*}\Sigma$ as a function of $\cos\theta$, specifically for the $N^{*}(2080)$ with $J^{P}=\frac{3}{2}^{-}$ at a center-of-mass energy $W=3.056$ GeV.
+3, we present the differential cross section for the reaction $\pi^{-}p\rightarrow K^{*}\Sigma$ as a function of $\cos\theta$, specifically for the $P_{s\bar{s}}$ with $J^{P}=\frac{3}{2}^{-}$ at a center-of-mass energy $W=3.056$ GeV.
 The analysis reveals that the $t$-channel contribution dominates at forward angles, while the $u$-channel contribution is small and negligible.
 In Fig.
-4, we show the differential cross section for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction as a function of center-of-mass energy, with $\cos\theta=1$, for the same $N^{*}(2080)$ with $J^{P}=\frac{3}{2}^{-}$.
-Since these measurements are taken at forward angles, the distinct shapes of the curves resulting from the contributions of the three channels allow us to effectively isolate the contributions from the $t$-channel and $u$-channel, while providing a clear indication of the $s$-channel $N^{*}(2080)(\frac{3}{2}^{-})$.
-This approach is ideal for studying the contribution of the $N^{*}(2080)(\frac{3}{2}^{-})$ through differential cross sections at forward angles.
+4, we show the differential cross section for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction as a function of center-of-mass energy, with $\cos\theta=1$, for the same $P_{s\bar{s}}$ with $J^{P}=\frac{3}{2}^{-}$.
+Since these measurements are taken at forward angles, the distinct shapes of the curves resulting from the contributions of the three channels allow us to effectively isolate the contributions from the $t$-channel and $u$-channel, while providing a clear indication of the $s$-channel $P_{s\bar{s}}[3/2^{-}]$.
+This approach is ideal for studying the contribution of the $P_{s\bar{s}}[3/2^{-}]$ through differential cross sections at forward angles.
 In Fig.
@@
 Conversely, as energy increases, the contribution of the $t$-channel at forward angles becomes more significant, while the $u$-channel contribution remains consistently small.
-These findings suggest that measuring the differential cross section of the reaction $\pi^{-}p\rightarrow K^{*}\Sigma$ near the $K^{*}\Sigma$ threshold will provide valuable information about the $N^{*}(2080)(\frac{3}{2}^{-})$ resonance.
+These findings suggest that measuring the differential cross section of the reaction $\pi^{-}p\rightarrow K^{*}\Sigma$ near the $K^{*}\Sigma$ threshold will provide valuable information about the $P_{s\bar{s}}[3/2^{-}]$ resonance.
 Figure 4: The differential cross-section $d\sigma/d\cos\theta$ for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction varies with different c.m.
@@
 2.
-The $t$-distribution for the reaction $\pi^{-}p\rightarrow K^{*}\Sigma$ with the $N^{*}(2080)$ and $J^{P}=\frac{3}{2}^{-}$ at $W=2.88$ GeV is shown in Fig.
+The $t$-distribution for the reaction $\pi^{-}p\rightarrow K^{*}\Sigma$ with the $P_{s\bar{s}}$ and $J^{P}=\frac{3}{2}^{-}$ at $W=2.88$ GeV is shown in Fig.
 6.
@@
 Figure 7: The $t$-distribution for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction at different c.m. energies $W$ = 2.2 GeV, 2.3 GeV, 2.4 GeV and 2.5 GeV. Here, the notation is the same as in Fig. 2.
-III.2 Production of the $N^{*}(2080)$ with $J^{P}$ = $\frac{1}{2}^{-}$
-Similar to the analysis of $N^{*}(2080)$ with $J^{P}$ = $\frac{3}{2}^{-}$, we also calculate the cross section for the reaction $\pi^{-}p\rightarrow K^{*}\Sigma$ with $N^{*}(2080)$ having $J^{P}$ = $\frac{1}{2}^{-}$ as detailed in Figs.
+III.2 Production of the $P_{s\bar{s}}$ with $J^{P}$ = $\frac{1}{2}^{-}$
+Similar to the analysis of $P_{s\bar{s}}$ with $J^{P}$ = $\frac{3}{2}^{-}$, we also calculate the cross section for the reaction $\pi^{-}p\rightarrow K^{*}\Sigma$ with $P_{s\bar{s}}$ having $J^{P}$ = $\frac{1}{2}^{-}$ as detailed in Figs.
 8, 9 and 10.
@@
 The fitting parameters are shown in Table 2.
-The fit excluding the $N^{*}(2080)(\frac{1}{2}^{-})$ contribution yields a $\chi^{2}/d.o.f.$ = 6.19, whereas including this contribution results in $\chi^{2}/d.o.f.$ = 2.20.
-For the $N^{*}(2080)$ with $J^{P}$ = $\frac{3}{2}^{-}$, the $\chi^{2}/d.o.f.$ = 1.52, indicating a better fit to the experimental data for the $N^{*}(2080)(\frac{3}{2}^{-})$.
+The fit excluding the $P_{s\bar{s}}[1/2^{-}]$ contribution yields a $\chi^{2}/d.o.f.$ = 6.19, whereas including this contribution results in $\chi^{2}/d.o.f.$ = 2.20.
+For the $P_{s\bar{s}}$ with $J^{P}$ = $\frac{3}{2}^{-}$, the $\chi^{2}/d.o.f.$ = 1.52, indicating a better fit to the experimental data for the $P_{s\bar{s}}[3/2^{-}]$.
 Fig.
@@
 energies at $\cos\theta$ = 1 is presented.
-This provides an ideal means to assess the contribution of the $N^{*}(2080)(\frac{1}{2}^{-})$ through forward-angle differential cross sections.
+This provides an ideal means to assess the contribution of the $P_{s\bar{s}}[1/2^{-}]$ through forward-angle differential cross sections.
 Fig.
@@
 With increasing energy, the $t$-channel contribution at forward angles grows more significant, while the $u$-channel contribution remains small.
-This behavior is similar to that observed for the $N^{*}(2080)(\frac{3}{2}^{-})$.
+This behavior is similar to that observed for the $P_{s\bar{s}}[3/2^{-}]$.
 Fig.
@@
 energies.
-These results may aid us in distinguishing the production mechanisms for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction with the $N^{*}(2080)$ having $J^{P}$ = $\frac{1}{2}^{-}$.
+These results may aid us in distinguishing the production mechanisms for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction with the $P_{s\bar{s}}$ having $J^{P}$ = $\frac{1}{2}^{-}$.
 Table 2: Fitted values of the free parameters with all experimental data in Refs.
-CERN-CollegedeFrance-Madrid-Stockholm:1980ysu ; CERN-CollegedeFrance-Madrid-Stockholm:1980mch ; Crennell:1972km ; Abramovich:1972rq ; Dahl:1967pg ; Miller:1965 for the case of the $N^{*}(2080)$ with $J^{P}=\frac{1}{2}^{-}$.
-with $N^{*}(2080)(\frac{1}{2}^{-})$ without $N^{*}(2080)(\frac{1}{2}^{-})$
+CERN-CollegedeFrance-Madrid-Stockholm:1980ysu ; CERN-CollegedeFrance-Madrid-Stockholm:1980mch ; Crennell:1972km ; Abramovich:1972rq ; Dahl:1967pg ; Miller:1965 for the case of the $P_{s\bar{s}}$ with $J^{P}=\frac{1}{2}^{-}$.
+with $P_{s\bar{s}}[1/2^{-}]$ without $P_{s\bar{s}}[1/2^{-}]$
 $\Lambda_{K}$ 1.81 $\pm$0.02 1.90$\pm$0.01
@@
 $\Lambda_{s}$ 1.61$\pm$0.02
-$g_{\pi N{N^{*}}}^{1/2^{-}}$ 0.20$\pm$0.01
+$g_{\pi N{P^{*}}}^{1/2^{-}}$ 0.20$\pm$0.01
 $\chi^{2}/d.o.f.$ 2.20 6.19
@@
 The solid (purple), dashed-dotted (blue), dashed (red), dash–double dotted (green) lines are for the full model, the $s$-channel, the $t$-channel and the $u$-channel, respectively.
-Here, the spin-parity quantum number of the $N^{*}(2080)$ is $\frac{1}{2}^{-}$.
+Here, the spin-parity quantum number of the $P_{s\bar{s}}$ is $\frac{1}{2}^{-}$.
 Figure 9: The differential cross-section $d\sigma/d\cos\theta$ for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction as a function of $\cos\theta$.
@@
 Figure 13: The $t$-distribution for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction at different c.m. energies $W$ = 2.2 GeV, 2.3 GeV, 2.4 GeV and 2.5 GeV. Here, the notation is the same as in Fig. 8.
-IV constituent counting rule in hard exclusive reactions
-We introduce the constituent counting rule for hard exclusive reactions by presenting the expressions for the cross section and matrix element, and by explaining a typical hard gluon exchange process that contributes to such reactions.
-For a large-angle exclusive scattering process $a+b\rightarrow c+d$, the reaction cross section is given by:
-$\frac{d\sigma_{ab\to cd}}{dt}\simeq\frac{1}{16\pi s^{2}}\overline{\sum_{\mathrm{pol}}}|M_{ab\to cd}|^{2},$ (30)
-In this context, the approximation implies that particle masses are neglected. The Mandelstam variables $s$ and $t$ are defined as follows:
-[…]
-where $p_{i}$ ($i=a,b,c,d$) represents the momentum of hadron $i$, while $p_{\mathrm{c.m.}}$ and $\theta_{\mathrm{c.m.}}$ denote the momentum and scattering angle in the center of mass (c.m.) frame, respectively.
-Eq.
-(30) includes an average over the initial spins and a sum over the final spins.
-The matrix element $M_{ab\rightarrow cd}$ can be expressed in a factorized form when there is a large momentum transfer Lepage:1980fj ; Mueller:1981sg ; Lepage:1979zb ; Efremov:1979qk ; Guo:2016fqg :
-[…]
-Here, $[x]$ represents the set of light-cone momentum fractions of partons within a hadron, where $x_{i}=p^{+}_{i}/p^{+}$, with $p_{i}$ and $p$ denoting the momenta of the $i$th parton and the hadron, respectively.
-The light-cone component is defined as $p^{+}=(p^{0}+p^{3})/\sqrt{2}$, considering the third axis as the longitudinal direction.
-For a nucleon, two independent variables are required to describe the distribution amplitude, while accounting for the constraint of momentum conservation, $x_{1}+x_{2}+x_{3}=1$.
-Thus, we have $[x]=x_{1},x_{2}$ and $[dx]=dx_{1}dx_{2}$ in Eq.
-(32).
-Conversely, only one variable, $x$, is needed for mesons such as pions and kaons.
-The variable $Q^{2}$ represents the hard scale of the reaction, approximately given by $Q^{2}\simeq s$ for large-angle elastic exclusive scattering.
-In Eq.
-(32), the dependencies on the renormalization and factorization scales are suppressed: the former is governed by the renormalization group equation for the coupling constant, and the latter by the Efremov-Radyushkin–Brodsky–Lepage (ERBL) evolution equations for the distribution amplitudes Lepage:1979zb ; Efremov:1979qk .
-These scales are typically taken to be on the order of $Q^{2}$ to minimize large radiative corrections.
-Consequently, the $Q^{2}$ dependencies are logarithmic and do not significantly affect the scaling behavior of the matrix element.
-Originally, the constituent counting rule was proposed based on dimensional counting Brodsky:1973kr ; Brodsky:1974vy ; Matveev:1973ra .
-It was subsequently explored within the framework of perturbative QCD, incorporating hard scattering processes Farrar:1979aw ; Efremov:1978rn ; Duncan:1979hi ; Duncan:1980qd ; Lepage:1980fj .
-In this section, we derive the scaling rule by counting mass dimensions.
-Given that the state vector of a hadron is normalized as $\langle h(p^{\prime})|h(p)\rangle=2p^{0}(2\pi)^{3}\delta^{(3)}(\vec{p}-\vec{p}^{\prime})$, we find that its mass dimension is $[1/M]$.
-If a hadron is composed of $n_{l}$ elementary constituents, its state vector can be expressed as
-$|l\rangle=\sqrt{N_{l}}|n_{l}\rangle,\quad\left[\sqrt{N_{l}}\right]=\left[M^{n_{l}-1}\right].$ (33)
-The second equation indicates that the normalization factor $\sqrt{N_{l}}$ carries the mass dimension $\left[M^{n_{l}-1}\right]$ when the state vector of each constituent has the mass dimension $[1/M]$.
-In practice, this normalization factor corresponds to the “decay constant,” which is independent of the hard momentum scale Kawamura:2013iia .
-We now consider the mass dimensions of the matrix element in Eq.
-(30) to derive the counting rule for exclusive cross sections.
-The scattering matrix $S$ is connected to the transition matrix $T$ by the relation $S=\bm{1}+i(2\pi)^{4}\delta^{(4)}(p_{f}-p_{i})T$, indicating that the mass dimension of $T$ is $[T]=[M^{4}]$.
-The matrix element $M_{ab\rightarrow cd}$ is then expressed in terms of $T$ as follows:
-$M_{ab\to cd}=\langle cd|T|ab\rangle=\sqrt{N_{a}N_{b}N_{c}N_{d}}\langle n_{c}n_{d}|T|n_{a}n_{b}\rangle.$ (34)
-Since the normalization factors $N_{i}(i=a,b,c,d)$ are represented by soft constants, we analyze the matrix element $\hat{M}_{ab\rightarrow cd}$ by omitting these factors.
-Consequently, the remaining hard component should be expressed in terms of the variables $s$ and $t$:
-[…]
-Given the dimensions $[T]=[M^{4}]$ and $[|n_{i}\rangle]=[1/M^{n_{i}}]$, the dimension of the matrix element is given by
-$\begin{bmatrix}\hat{M}_{ab\to cd}\end{bmatrix}=\begin{bmatrix}\langle n_{c}n_{d}|T|n_{a}n_{b}\rangle\end{bmatrix}=\begin{bmatrix}M^{4-n}\end{bmatrix},$ (36)
-where $n\equiv n_{a}+n_{b}+n_{c}+n_{d}$.
-In a large-angle exclusive reaction, the variable $s$ can be chosen as the sole hard scale.
-Thus, considering the mass dimension, the matrix element is expressed as:
-$\hat{M}_{ab\to cd}=\hat{F}_{ab\to cd}(s,t)=s^{(4-n)/2}F_{ab\to cd}(t/s),$ (37)
-where $F_{ab\to cd}(t/s)$ is a dimensionless function of scattering angle, with $-2t/s=1-\cos\theta_{c.m.}$ from Eq.
-(31).
-Using Eqs.
-(30), (35), and (37), we derive the constituent-counting expression for the cross section:
+III.3 A further test by the constituent counting rule
+In the following, we adopt the approach of the constituent counting rule to further test our theoretical calculation around the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction.
+Generally, for a large-angle exclusive scattering process $a+b\rightarrow c+d$, the reaction cross section can be expressed with the constituent counting rule (see Refs.
+Lepage:1980fj ; Mueller:1981sg ; Lepage:1979zb ; Efremov:1979qk ; Brodsky:1973kr ; Brodsky:1974vy ; Matveev:1973ra ; Efremov:1978rn ; Duncan:1979hi ; Duncan:1980qd ; Lepage:1980fj ; Farrar:1979aw ; Kawamura:2013iia ; Anderson:1976ph ; JeffersonLabHallA:2002vjd ; JeffersonLabHallA:2004gxp ; Baller:1988tj ; White:1994tj for more details):
 $\frac{d\sigma_{ab\to cd}}{dt}=\frac{1}{s^{n-2}}f_{ab\to cd}(t/s),$ (30)
-where $f(t/s)$ accounts for the scattering angle and is multiplied by the normalization factors.
-The mass dimensions of $f(t/s)$ are given by $[f(t/s)]=[N_{a}N_{b}N_{c}N_{d}|F_{ab\to cd}|^{2}]=[M^{2n-8}]$, so the overall mass dimension of Eq.
-(38) is $[1/M^{4}]$.
-This derivation illustrates the counting rule based on mass dimensions, referred to as the “constituent counting rule” since it quantifies the number of constituents involved in the reaction.
-Experimentally, the constituent counting rule was first applied in Ref.
-Anderson:1976ph to reactions such as $\gamma p\rightarrow\pi^{+}n$, $\pi^{0}p$, $\pi^{-}\Delta^{++}$, $\rho^{0}p$, $K^{+}\Lambda$, and $K^{+}\Sigma^{0}$, suggesting $n=3$ for baryons and $n=2$ for mesons.
-The JLab (Thomas Jefferson National Accelerator Facility) Hall-A and E94-104 collaborations subsequently verified that experimental data are consistent with $n_{\gamma}=1$, $n_{p}=n_{n}=3$, and $n_{\pi^{+}}=2$ by analyzing the $\gamma p\rightarrow\pi^{+}n$ reaction JeffersonLabHallA:2002vjd ; JeffersonLabHallA:2004gxp .
-Furthermore, several experiments have confirmed the scaling behavior of the exclusive cross section predicted by the constituent counting rule Baller:1988tj ; White:1994tj .
-Table 3: The cross sections of $\pi^{-}p\rightarrow K^{*}\Sigma$ at $\theta_{\mathrm{c.m.}}=90^{\circ}$ for the case of the $N^{*}(2080)$ with $J^{P}$ = $\frac{3}{2}^{-}$.
+where $n\equiv n_{a}+n_{b}+n_{c}+n_{d}$, while $f(t/s)$ accounts for the scattering angle and is multiplied by the normalization factors.
+Table 3: The cross sections of $\pi^{-}p\rightarrow K^{*}\Sigma$ at $\theta_{\mathrm{c.m.}}=90^{\circ}$ for the case of the $P_{s\bar{s}}$ with $J^{P}$ = $\frac{3}{2}^{-}$.
 $\sqrt{s}$ (GeV) $t$ $(\mathrm{GeV}^{2})$ $\frac{d\sigma}{dt}$ $(\mu b/{\mathrm{GeV}}^{2})$
@@
 2.80 2.393 2.250$\pm$0.970
-Table 4: The cross sections of $\pi^{-}p\rightarrow K^{*}\Sigma$ at $\theta_{\mathrm{c.m.}}=90^{\circ}$ for the case of the $N^{*}(2080)$ with $J^{P}$ = $\frac{1}{2}^{-}$.
+Table 4: The cross sections of $\pi^{-}p\rightarrow K^{*}\Sigma$ at $\theta_{\mathrm{c.m.}}=90^{\circ}$ for the case of the $P_{s\bar{s}}$ with $J^{P}$ = $\frac{1}{2}^{-}$.
 $\sqrt{s}$ (GeV) $t$ (GeV2) $\frac{d\sigma}{dt}$ $(\mu b/{\mathrm{GeV}}^{2})$
@@
 2.80 2.393 3.122$\pm$0.509
-In this study, we test the constituent counting rule using the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction.
-According to the rule, the total number of constituents is $n=2+3+2+3=10$.
-We calculate the cross sections for the $\pi^{-}p\rightarrow K^{*}\Sigma$ process at $\theta_{\mathrm{c.m.}}=90^{\circ}$, considering the cases where the $N^{*}(2080)$ has $J^{P}=\frac{3}{2}^{-}$ and $J^{P}=\frac{1}{2}^{-}$.
+According to the rule, in the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction, the total number of constituents is $n=2+3+2+3=10$.
+We calculate the cross sections for the $\pi^{-}p\rightarrow K^{*}\Sigma$ process at $\theta_{\mathrm{c.m.}}=90^{\circ}$, considering the cases where the $P_{s\bar{s}}$ has $J^{P}=\frac{3}{2}^{-}$ and $J^{P}=\frac{1}{2}^{-}$.
 The results are presented in Table 3 and Table 4, respectively.
 By fitting the experimental data at $\sqrt{s}=2.88$ GeV and the numerical results shown in Table 3 and Table 4 to the expression $d\sigma/dt=(\text{constant})\times s^{2-n}$, the scaling factors obtained are presented in Table 5.
-We find that for both $N^{*}(2080)(\frac{3}{2}^{-})$ and $N^{*}(2080)(\frac{1}{2}^{-})$, the fitted values of $n$ are very close to 10, confirming the accuracy of the constituent counting rule.
+We find that the fitted values of $n$ are very close to 10, confirming the accuracy of the theoretical prediction for the cross section,.
 Fig.
 14 and Fig.
-15 depict the differential cross section $d\sigma/dt$ of the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction at a large meson production angle ($\theta=90^{\circ}$ in the center of mass frame) as a function of $\sqrt{s}$ for the $N^{*}(2080)(\frac{3}{2}^{-})$ and $N^{*}(2080)(\frac{1}{2}^{-})$, respectively.
+15 depict the differential cross section $d\sigma/dt$ of the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction at a large meson production angle ($\theta=90^{\circ}$ in the center of mass frame) as a function of $\sqrt{s}$ for the $P_{s\bar{s}}[3/2^{-}]$ and $P_{s\bar{s}}[1/2^{-}]$, respectively.
+Both figures clearly demonstrate that as energy increases, the theoretically predicted differential cross section aligns more closely with the constituent counting rules.
+This occurs because these rules are generally more applicable in high-energy, large-momentum transfer regions.
+In this high-energy regime, which is distant from the mass threshold of $P_{s\bar{s}}$, the contribution of $P_{s\bar{s}}$ production to the differential cross section becomes minimal.
+Consequently, this region does not allow the constituent counting rules to reflect the properties of the intermediate exchange particle $P_{s\bar{s}}$.
+Introducing the constituent counting rules serves to verify the theoretical cross-section calculations presented in this work, further supporting the reliability of our theoretical model.
 Table 5: Fitted values of the scaling factor $n$ by using the $\chi^{2}$ fitting algorithm.
-$N^{*}(2080)(\frac{3}{2}^{-})$ $N^{*}(2080)(\frac{1}{2}^{-})$
+$P_{s\bar{s}}[3/2^{-}]$ $P_{s\bar{s}}[1/2^{-}]$
 $n$ 10.01$\pm$0.01 10.40$\pm$0.04
@@
 The band stands for the error bar of the fitting parameters in Table 5.
-In this work, the constituent counting rule of perturbative QCD could be utilized to determine the internal quark configuration of exotic-hadron candidates in exclusive production processesKawamura:2013iia .
-In the exclusive process $a+b\rightarrow c+d$ with large-momentum transfer, quarks should share large momenta so that they can stick together to form a hadron by exchanging hard gluons.
-Since the cross section of the $a+b\rightarrow c+d$ exclusive reaction should scale like $d\sigma/dt\sim s^{2-n}f(\theta_{\mathrm{c.m.}})$ with $n=n_{a}+n_{b}+n_{c}+n_{d}$, where $n_{i}(i=a,b,c,d)$ is the number of constituents in the particle $i$.
-Since the factor $n_{i}$ clearly indicates the internal configuration in the hadron, this scaling relation can be employed to find the internal configurations of exotic hadrons.
-The constituent counting rule provides useful guidance for us to further explore the properties of exotic hadrons.
-V Dalitz process and experimental feasibility
+IV Dalitz process and experimental feasibility
 Figure 16: Feynman diagrams for the $\pi^{-}p\rightarrow K^{+}\pi^{-}\Sigma^{0}$ reaction.
@@
 16.
-Our calculations show that the total cross section for $N^{*}(2080)(\frac{1}{2}^{-})$ is slightly larger than that for the $N^{*}(2080)(\frac{3}{2}^{-})$.
-In this study, we focus on the Dalitz process specifically for $N^{*}$ with $J^{P}=\frac{3}{2}^{-}$.
+Our calculations show that the total cross section for $P_{s\bar{s}}[1/2^{-}]$ is slightly larger than that for the $P_{s\bar{s}}[3/2^{-}]$.
+In this study, we focus on the Dalitz process specifically for $P_{s\bar{s}}$ with $J^{P}=\frac{3}{2}^{-}$.
 Generally, the invariant mass spectrum of the Dalitz process is defined based on the two-body process Kim:2017nxg ; Wang:2024qnk
@@
 The minimum ratio reaches 29.67% at center-of-mass energies between 2.56 and 2.88 GeV, indicating that the total cross section and event number for the $\pi^{-}p\rightarrow K^{*}\Sigma$ process reconstructed via $\pi^{-}p\rightarrow K^{+}\pi^{-}\Sigma^{0}$ are sufficient for experimental measurements.
-Additionally, the ratio can reach up to 76.61% at a center-of-mass energy of 2.56 GeV, with the $s$-channel $N^{*}(2080)$ exchange contributing 73.19%, approximately 30.28 $\mu\text{b}$.
-These results suggest that a significant number of the $N^{*}(2080)$ can be produced in the $\pi^{-}p\rightarrow K^{*}\Sigma\rightarrow K^{+}\pi^{-}\Sigma^{0}$ process, providing strong support for future experimental efforts.
+Additionally, the ratio can reach up to 76.61% at a center-of-mass energy of 2.56 GeV, with the $s$-channel $P_{s\bar{s}}$ exchange contributing 73.19%, approximately 30.28 $\mu\text{b}$.
+These results suggest that a significant number of the $P_{s\bar{s}}$ can be produced in the $\pi^{-}p\rightarrow K^{*}\Sigma\rightarrow K^{+}\pi^{-}\Sigma^{0}$ process, providing strong support for future experimental efforts.
 Table 6: The total cross section of $\pi^{-}p\rightarrow K^{+}\pi^{-}\Sigma^{0}$ and $\pi^{-}p\rightarrow K^{*}\Sigma\rightarrow K^{+}\pi^{-}\Sigma^{0}$.
@@
 As a leading hadron experimental facility, J-PARC currently provides the highest-intensity meson beams (both $K$ and $\pi$ mesons) in the several GeV energy range, making it an outstanding platform for particle and nuclear physics research Aoki:2021cqa .
-Theoretical studies and experimental evidence suggest that certain light-flavored baryons may contain a significant strange quark pair ($s\bar{s}$) component, potentially forming a hidden-strange pentaquark state, such as the $N^{*}(2080)$.
+Theoretical studies and experimental evidence suggest that certain light-flavored baryons may contain a significant strange quark pair ($s\bar{s}$) component, potentially forming a hidden-strange pentaquark state, such as the $P_{s\bar{s}}$.
 However, the existence of a hidden-strange pentaquark state has not yet been confirmed experimentally, underscoring the need for collaborative efforts between experimental and theoretical research.
-If J-PARC can measure the $\pi^{-}p\rightarrow K^{*}\Sigma$ process, the findings could aid in the discovery of the hidden-strange pentaquark state $N^{*}(2080)$ and provide valuable insights into quark interactions within hadrons, particularly the internal dynamics of diquarks.
+If J-PARC can measure the $\pi^{-}p\rightarrow K^{*}\Sigma$ process, the findings could aid in the discovery of the hidden-strange pentaquark state $P_{s\bar{s}}$ and provide valuable insights into quark interactions within hadrons, particularly the internal dynamics of diquarks.
 Additionally, HIAF offers high-intensity proton and ion beam currents across a wide energy range.
 Upon completion, HIAF will deliver the highest peak-intensity medium and low-energy heavy-ion beam currents compared to similar facilities worldwide HIAF .
-If a meson beam experimental device is developed at HIAF in the future, it could enable high-precision measurements of the $\pi^{-}p\rightarrow K^{*}\Sigma$ process, further supporting the search for the $N^{*}(2080)$ wl:2024ext .
-VI Summary and prospect
-In recent years, the $N^{*}(2080)$ has been considered the hidden-strange partner of the hidden-charm molecular pentaquark $P_{c}(4457)$, as its mass lies just below the $K^{*}\Sigma$ threshold He:2017aps ; Ben:2023uev ; Xie:2010yk ; Lin:2018kcc .
-In this study, we investigate the production of the nucleon excited state $N^{*}(2080)$ through the $\pi^{-}p$ scattering process using the effective Lagrangian approach.
-Specifically, we consider $K$ and $K^{*}$ meson exchanges in the $t$-channel, $\Sigma$ exchange in the $u$-channel, and the contribution of the $N^{*}(2080)$ in the $s$-channel.
-By fitting the total and differential cross sections of the $\pi^{-}p\rightarrow K^{*}\Sigma$ process, we demonstrate that the $N^{*}(2080)$ exchange provides the dominant contribution to the total cross section near the threshold energy region, offering potential insights into detecting the $N^{*}(2080)$ through the $\pi^{-}p$ scattering process.
-However, the lack of experimental data near the threshold imposes limitations on determining the properties of the $N^{*}(2080)$ via this process.
+If a meson beam experimental device is developed at HIAF in the future, it could enable high-precision measurements of the $\pi^{-}p\rightarrow K^{*}\Sigma$ process, further supporting the search for the $P_{s\bar{s}}$ wl:2024ext .
+V Summary and prospect
+In recent years, the $P_{s\bar{s}}$ has been considered the hidden-strange partner of the hidden-charm molecular pentaquark $P_{c}(4457)$, as its mass lies just below the $K^{*}\Sigma$ threshold He:2017aps ; Ben:2023uev ; Xie:2010yk ; Lin:2018kcc .
+In this study, we investigate the production of the hidden-strange molecular-type pentaquarks $P_{s\bar{s}}$ through the $\pi^{-}p$ scattering process using the effective Lagrangian approach.
+Specifically, we consider $K$ and $K^{*}$ meson exchanges in the $t$-channel, $\Sigma$ exchange in the $u$-channel, and the contribution of the $P_{s\bar{s}}$ in the $s$-channel.
+By fitting the total and differential cross sections of the $\pi^{-}p\rightarrow K^{*}\Sigma$ process, we demonstrate that the $P_{s\bar{s}}$ exchange provides the dominant contribution to the total cross section near the threshold energy region, offering potential insights into detecting the $P_{s\bar{s}}$ through the $\pi^{-}p$ scattering process.
+However, the lack of experimental data near the threshold imposes limitations on determining the properties of the $P_{s\bar{s}}$ via this process.
 Given the significant branching ratio of $K^{*}$ decay into $K^{+}\pi^{-}$, we calculate the Dalitz process for $\pi^{-}p\rightarrow K^{*}\Sigma\rightarrow K^{+}\pi^{-}\Sigma^{0}$ and compare it with experimental data for $\pi^{-}p\rightarrow K^{+}\pi^{-}\Sigma^{0}$.
-The results suggest that it is feasible to observe the $N^{*}(2080)$ through the $\pi^{-}p\rightarrow K^{*}\Sigma$ process in experiments.
-Notably, our calculations consider two scenarios for the $J^{P}$ quantum number of the $N^{*}(2080)$, either $\frac{1}{2}^{-}$ or $\frac{3}{2}^{-}$.
-A significant source of uncertainty in our model arises from the absence of experimental measurements of the partial widths of the $N^{*}(2080)$ decays to $\pi N$, leading us to treat the coupling constants $g_{\pi NN^{*}}^{3/2^{-}}$ and $g_{\pi NN^{*}}^{1/2^{-}}$ as free parameters determined by fitting.
-In conclusion, further experimental measurements and theoretical investigations into the internal properties of $N^{*}$ states are highly desirable.
-Additionally, the application of the constituent counting rule in hard exclusive reactions is discussed, including a demonstration of the rule by deriving expressions for the reaction cross-section and matrix element, as well as explaining the contribution of typical hard gluon exchange processes to exclusive reactions.
-By calculating the mass dimension, we derive the constituent counting expression for the reaction cross-section.
-In the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction, the predictions of this rule align with experimental data, further validating its accuracy.
+The results suggest that it is feasible to observe the $P_{s\bar{s}}$ through the $\pi^{-}p\rightarrow K^{*}\Sigma$ process in experiments.
+Notably, our calculations consider two scenarios for the $J^{P}$ quantum number of the $P_{s\bar{s}}$, either $\frac{1}{2}^{-}$ or $\frac{3}{2}^{-}$.
+A major source of uncertainty in our model stems from the lack of experimental data on the partial widths of $P_{s\bar{s}}$ decays to $\pi N$.
+This necessitates treating the coupling constants $g_{\pi NP^{*}}^{3/2^{-}}$ and $g_{\pi NP^{*}}^{1/2^{-}}$ as free parameters, determined through fitting.
+In these fitting calculations, both the coupling constants and the cutoff parameters in the form factors are set as free parameters, which is an inherent aspect of this theoretical approach.
+Consequently, while we can calculate and analyze the production mechanism and cross section of the $P_{s\bar{s}}$ state via the $\pi^{-}p\rightarrow K^{*}\Sigma$ process, we cannot further deduce the internal structural properties of the $P_{s\bar{s}}$ state.
+Additionally, the application of the constituent counting rule in hard exclusive reactions is discussed.
+For the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction, our theoretical cross section predictions align well with the constituent counting rules, providing a strong validation of the overall cross-section calculation.
 Currently, high-precision measurements of the meson-nucleus scattering process can be conducted by experiments at J-PARC Aoki:2021cqa , AMBER Adams:2018pwt , and future facilities such as HIKE HIKE:2023ext and HIAF HIAF .
 We recommend that these experiments perform precise measurements of the $\pi^{-}p\rightarrow K^{*}\Sigma$ process, particularly the differential cross section at the forward angle ($\theta=0^{\circ}$) and the $t$-distribution cross section at large momentum transfer ($\theta=90^{\circ}$).
-This is crucial for elucidating the mechanism of the $\pi^{-}p\rightarrow K^{*}\Sigma$ process and determining the contribution of the $N^{*}$ state.
+This is crucial for elucidating the mechanism of the $\pi^{-}p\rightarrow K^{*}\Sigma$ process and determining the contribution of the $P_{s\bar{s}}$ state.
 To date, pentaquark-like states containing strange quarks have not been experimentally confirmed, necessitating further clarification through collaborative experimental and theoretical efforts.
@@
 We will also collaborate closely with experimentalists to conduct further in-depth research in this area.
-VII Acknowledgments
+VI Acknowledgments
 This work is supported by the National Natural Science Foundation of China under Grants No.